### v2

# The directed landscape is a black noise

Abstract.
We show that the directed landscape is a black noise in the sense of Tsirelson and Vershik. As a corollary, we show that for any microscopic system in which the height profile converges in law to the directed landscape, the driving noise is asymptotically independent of the height profile. This decoupling result provides one answer to the question of what happens to the driving noise in the limit under the KPZ scaling, and illustrates a type of noise sensitivity for systems in the KPZ universality class. Such decoupling and sensitivity phenomena are not present in the intermediate-disorder or weak-asymmetry regime, thus illustrating a contrast from the weak KPZ scaling regime. Along the way, we prove a strong mixing property for the directed landscape on a bounded time interval under spatial shifts, with a mixing rate $\alpha(k)\leq Ce^{-dk^{3}}$ for some $C,d>0$.
1. Introduction
1.1. Preface
Over the past several decades, mathematicians and physicists who study interacting particle systems, integrable systems, and the Kardar-Parisi-Zhang (KPZ) equation have discovered connections between the large-scale behavior of a wide-ranging class of models. It is known that the height functions of asymmetric exclusion processes and stochastic vertex models exhibit universal scaling behavior in $1+1$ dimensions. A common feature in these models is the appearance of $1:2:3$ scaling exponents and Tracy-Widom limit laws. Similar results are expected for the Eden model and last passage percolation models with arbitrary weight distributions, as well as Burgers-type stochastic PDEs and ballistic deposition models (e.g. Tetris). This phenomenon is called KPZ universality [Cor12, Qua11].
In recent years, there has been a great deal of work involved in constructing and understanding the universal limit objects in the KPZ universality class. Matetski, Quastel, and Remenik [MQR21] constructed the KPZ fixed point, the scale-invariant limiting Markov process (conjecturally) of all models in the KPZ universality class. Dauvergne, Ortmann and Virág [DOV22] then constructed the directed landscape, a generalization which encodes the natural joint couplings of all KPZ fixed points started from all possible initial data simultaneously. The directed landscape is a four-parameter field, denoted in this paper as $\mathcal{L}_{s,t}(x,y)$, see 1.8.
It is a longstanding project of mathematical physicists to have a better understanding of this phenomenon of universality. It is conjectured that any model for which it is possible to associate a height function $h(t,x)$ with the basic properties of local dynamics, a smoothing mechanism, lateral growth, and space-time forcing with rapid correlation decay, will converge to the KPZ fixed point under $1:2:3$ scaling $\epsilon h(\epsilon^{-3}t,\epsilon^{-2}x)-C_{\epsilon}t$, see for example the introduction of [MQR21]. More generally, the height mechanisms started simultaneously from several distinct initial data should converge to KPZ fixed points coupled through the directed landscape. Models which satisfy these conditions, and which have been proved or conjectured to have the KPZ fixed point and directed landscape as a scaling limit, are said to belong to the KPZ universality class.
The directed landscape (1.8), which is the main object in this paper, was constructed by Dauvergne, Ortmann, and Virág [DOV22] as the scaling limit of the last passage metric between points on a Brownian line ensemble (Brownian LPP). It is a random four-parameter continuous function $\mathcal{L}:\mathbb{R}^{4}_{\uparrow}\to\mathbb{R}$, with $\mathbb{R}^{4}_{\uparrow}:=\{(x,s;y,t)\in\mathbb{R}^{4}|s<t\},$ which may be viewed as a metric between space-time points $(s,x)$ and $(t,y)$ that satisfies a reverse triangle inequality, $\mathcal{L}_{s,r}(x,z)\geq\mathcal{L}_{s,t}(x,y)+\mathcal{L}_{t,r}(y,z).$ More precisely the following “metric composition” law holds:
$\mathcal{L}_{s,r}(x,z)$ $=\max_{y\in\mathbb{R}}\left(\mathcal{L}_{s,t}(x,y)+\mathcal{L}_{t,r}(y,z)\right),$ for $\text{for all }(x,z)\in\mathbb{R}^{2},\;\;s<t<r.$ (1.1)
A series of recent works have demonstrated that the directed landscape is the scaling limit of many well-known models with simple microscopic dynamics. Works such as [DOV22, DNV23, DV21] demonstrated that it is the limit of the Brownian, Poissonian, exponential, and geometric last passage percolation, and Poisson and Seppäläinen-Johansson line ensembles. A work by Wu [Wu23] combined with a result of Quastel and Sarkar [QS20] demonstrates that the directed landscape is the scaling limit of the KPZ equation started from all possible initial data but coupled through the same realization of the driving noise. Most recently, a work by Aggarwal, Corwin, and Hegde [ACH24] demonstrated a similar convergence for ASEP and stochastic six vertex models. The KPZ fixed point of Matetski, Quastel, and Remenik [MQR21] is a marginal of the directed landscape by the variational formula $h_{t}(y)=\sup_{x\in\mathbb{R}}\left(h_{0}(x)+\mathcal{L}_{0,t}(x,y)\right)$, see [NQR20].
There are still many basic properties of the directed landscape that have yet to be explored. A natural question to ask is whether it has some simpler Gaussian structure.
Question 1.1.
Is the directed landscape the adapted strong solution of a white-noise driven SPDE?
This article was motivated by an interest in better understanding what happens to the driving noise of the KPZ equation under the directed landscape limit. Another natural question in this vein is the following.
Question 1.2.
What happens to the microscopic driving noise of systems like last passage percolation models, asymmetric exclusion processes, and the KPZ equation as we take the limit to the directed landscape? In particular, is any memory of this microscopic noise retained by the limiting directed landscape?
This paper aims to address both of the preceding questions. We study the directed landscape as an abstract noise in the sense of Tsirelson and Vershik [TV98], demonstrating that it is a black noise (1.10). Using this property, we show that the directed landscape is independent of any white noise field defined on the same filtered probability space (1.11). These results have a few corollaries. We conclude that the height function associated to any model which converges to the directed landscape will asymptotically decouple from the random environment (1.12), which answers 1.2. We can also answer 1.1 in the negative (1.13).
1.2. Background and Results
1.2.1. Black Noise
Black noise was first introduced by Tsirelson and Vershik [TV98], and subsequently explored at length in a survey by Tsirelson [Tsi04]. Building towards the definition of black noise, we begin by introducing the concept of a noise.
Definition 1.3 (Definition 3d(1), [Tsi04]).
A noise, denoted in this paper as a 4-tuple $(\Omega,(\mathcal{F}_{s,t})_{s<t},\mathbb{P},(\theta_{h})_{h})$, is a probability space $(\Omega,\mathcal{F},\mathbb{P})$ with sub-$\sigma$-fields $\mathcal{F}_{s,t}\subset\mathcal{F}$ ($-\infty<s<t<\infty$) and measurable maps $\theta_{h}:\Omega\to\Omega$ ($h\in\mathbb{R}$) such that the following properties are satisfied.
  1. (1)
    $\mathcal{F}_{s,t}$ and $\mathcal{F}_{t,u}$ are independent under the probability measure $P$ if $s<t<u$.
  2. (2)
    For all $r<s<t$, $\mathcal{F}_{r,s}$ and $\mathcal{F}_{s,t}$ together generate $\mathcal{F}_{r,t}$, i.e., $\mathcal{F}_{r,s}\vee\mathcal{F}_{s,t}=\mathcal{F}_{r,t}$.
  3. (3)
    $\mathcal{F}$ is generated by the union of all $\mathcal{F}_{s,t}.$
  4. (4)
    $\theta_{0}=\mathrm{Id}$, and $\theta_{h}\theta_{k}=\theta_{h+k}$ for all $h,k\in\mathbb{R}.$
  5. (5)
    If $A\in\mathcal{F}_{s,t}$ and $h\in\mathbb{R}$, then $\theta_{h}(A)\in\mathcal{F}_{s+h,t+h}$, and $P(\theta_{h}(A))=P(A).$
This definition of noise uses the language of stochastic flows, in the sense of Tsirelson [Tsi04]. We can relate this to other, possibly more familiar, definitions of noise. For example, if $B$ is a one-dimensional two-sided Brownian motion, the $\sigma$-algebra $\mathcal{F}_{s,t}$ would be generated by random variables of the form $B(u)-B(s)$ for $u\in[s,t]$, and $\theta_{h}B(u):=B(u+h)-B(h)$. More generally a Gaussian space-time white noise on ×$\mathbb{R}\times\mathbb{R}^{d}$, as discussed in Walsh’s book on SPDEs [Wal86], associates a Gaussian random variable to every Borel set, and defines a structure of a noise in a similar and straightforward fashion. These are both examples of white noise.
We can now give a definition of black noise. There are several equivalent definitions; Tsirelson [Tsi04] uses the notion of a stable $\sigma$-algebra to define white noise and black noise. The definition we use in this paper is more direct.
Definition 1.4.
Let $(\Omega,(\mathcal{F}_{s,t})_{s<t},\mathbb{P},(\theta_{h})_{h\in\mathbb{R}})$ be a noise. A random variable $F:\Omega\to\mathbb{R}$ is called linear if
$\mathbb{E}\left[F|\mathcal{F}_{s,t}\right]+\mathbb{E}\left[F|\mathcal{F}_{t,u}\right]=\mathbb{E}\left[F|\mathcal{F}_{s,u}\right]\;\;\;\;\;\;\;\;\;\mathrm{for\;\;all\;\;\;\;}s<t<u.$
We say that $(\Omega,(\mathcal{F}_{s,t})_{s<t},\mathbb{P},(\theta_{h})_{h\in\mathbb{R}})$ is a white noise if $\mathcal{F}$ is generated by linear random variables. Likewise, we say that $(\Omega,(\mathcal{F}_{s,t})_{s<t},\mathbb{P},(\theta_{h})_{h\in\mathbb{R}})$ is a black noise if 00 is the only linear random variable.
In general, white noise consists of Gaussian and Poissonian components, whereas black noise can be much more complex and admits no simple description. White noise and black noise do not form a dichotomy: there are noises that satisfy 1.3 which have both a “white noise part” and a “black noise part,” see [Tsi04, Section 6].
1.2.2. The KPZ Universality Class and the directed landscape
In this section, we define the directed landscape and several other important objects in the KPZ universality class. We begin by discussing the Airy process and Airy line ensemble.
Definition 1.5.
The Airy Line Ensemble is a collection of random functions $\left\{\mathcal{A}_{i}\right\}_{i\in\mathbb{N}}$ with law determined by the Airy kernel (1.2). For any finite set $I\subset\mathbb{R}$, we define the point process $\{(x,\mathcal{A}_{i}(x))|x\in I\}$ on $I\times\mathbb{R}$ as the determinantal point process [HKPV09, Section 4] with kernel $K(x_{1},t_{1};x_{2},t_{2})$ given by the extended Airy kernel
[…]
In this formula, $\mathrm{Ai}(\cdot)$ is the Airy function,
[…]
The parabolic Airy line ensemble is defined by $\mathcal{P}_{i}(x)=\mathcal{A}_{i}(x)-x^{2}$. We commonly refer to the top curve $\mathcal{P}_{1}(\cdot)$ as the parabolic Airy2 process.
The parabolic Airy line ensemble can also be defined as the unique line ensemble with a Brownian Gibbs resampling property and with the parabolic Airy2 process as its top line. Existence of such a line ensemble was shown by Corwin and Hammond [CH14] and uniqueness was shown by Dimitrov and Matetski [DM21].
For a metric space $X$, we let $C(X)$ denote the space of continuous functions $X\to\mathbb{R}$, equipped with the topology of uniform convergence on compacts. For the choices of $X$ used in this paper, the space $C(X)$ is always separable and metrizable, hence Polish. We can now introduce the Airy sheet, a two-variable random function that can be viewed as the universal limit of the last passage times for pairs of points of models in the KPZ class.
Definition 1.6.
The Airy sheet is a random continuous function $\mathcal{S}:\mathbb{R}^{2}\to\mathbb{R}$, i.e., a $C(\mathbb{R}^{2})$-valued random variable, such that the following properties hold:
  1. (1)
    $\mathcal{S}(\cdot,\cdot)$ has the same law as $\mathcal{S}(\cdot+t,\cdot+t)$ for all $t\in\mathbb{R}$, as $C(\mathbb{R}^{2})$-valued random variables.
  2. (2)
    $\mathcal{S}$ can be coupled with a parabolic Airy line ensemble so that $\mathcal{S}(0,\cdot)=\mathcal{P}_{1}(\cdot)$ and for all $(x,y,z)\in\mathbb{R}^{+}\times\mathbb{R}^{2}$,
    $\lim_{k\to\infty}\mathcal{P}\left[\left(-\sqrt{\frac{k}{2x}},k\right)\to\left(z,1\right)\right]-\mathcal{P}\left[\left(-\sqrt{\frac{k}{2x}},k\right)\to\left(y,1\right)\right]=\mathcal{S}(x,z)-\mathcal{S}(x,y),$
    where $\mathcal{P}\left[\left(a,k\right)\to\left(b,1\right)\right]$ is the last passage time between points $(a,k)$ and $(b,1)$ on the Airy line ensemble,
    $\mathcal{P}[(a,k)\to(b,1)]:=\sup_{a=t_{k}<...<t_{0}=b}\sum_{i=1}^{k}\left(\mathcal{P}_{i}(t_{i-1})-\mathcal{P}_{i}(t_{i})\right).$
Dauvergne, Ortmann, and Virág [DOV22] proved that the Airy sheet exists and that it is unique in law, in fact, it was later shown to be a deterministic functional of the Airy line ensemble [DV21]. With these definitions in hand, we can discuss the directed landscape. As explained in [DV21, Example 1.6], the directed landscape can be interpreted as a random directed spacetime metric of negative sign. We can think of the directed landscape as a universal limit of the metric defined by last passage times on a last passage model. We begin by defining the canonical probability space for the directed landscape.
Definition 1.7.
We define the canonical probability space of the directed landscape $(\Omega,\mathcal{F^{\mathcal{L}}})$. Let $\mathbb{R}^{4}_{\uparrow}:=\{(x,s;y,t)\in\mathbb{R}^{4}|s<t\}$. Define the underlying probability space $\Omega=C(\mathbb{R}^{4}_{\uparrow})$, equipped with its Borel sets from the topology of uniform convergence on compact sets. The canonical process on this space is the identity map $\mathcal{L}:C(\mathbb{R}^{4}_{\uparrow})\to C(\mathbb{R}^{4}_{\uparrow})$. The evaluation map at a point $(x,s,y,t)$ applied to $\mathcal{L}$ will be denoted in this paper as $\mathcal{L}_{s,t}(x,y)$. In this notation, we can define $\mathcal{F^{\mathcal{L}}}$ as the $\sigma$-algebra generated by
$\mathcal{F}^{\mathcal{L}}_{s,t}:=\sigma\left(\{\mathcal{L}_{a,b}(x,y):s\leq a<b\leq t,(x,y)\in\mathbb{R}^{2}\}\right),\;\;\;\;\;\;\;\;-\infty<s<t<\infty.$
Definition 1.8.
The directed landscape is a four-parameter field of random variables $\mathcal{L}_{s,t}(x,y)$, indexed by $-\infty<s<t<\infty$ and $x,y\in\mathbb{R}$, such that
  1. (1)
    $\mathcal{L}_{s,t}$ and $\mathcal{L}_{t,u}$ are independent for $s<t<u$.
  2. (2)
    $\mathcal{L}_{s,t}\circ\mathcal{L}_{t,u}=\mathcal{L}_{s,u}$ for all $s<t<u$ where the “metric composition” is defined for two functions $f,g:\mathbb{R}^{2}\to\mathbb{R}$ as
    $(f\circ g)(x,z)=\sup_{y\in\mathbb{R}}f(x,y)+g(y,z).$
  3. (3)
    For all $s<t$, $\mathcal{L}_{s,t}$ has the same law as $R_{t-s}\mathcal{L}_{0,1}$ where $R_{\epsilon}f(x,y):=\epsilon^{1/3}f(\epsilon^{-2/3}x,\epsilon^{-2/3}y).$
  4. (4)
    $\mathcal{L}_{0,1}$ is distributed as the Airy Sheet.
As an immediate consequence of this definition, the directed landscape defines a noise.
Lemma 1.9.
The directed landscape defines a noise $(\Omega,\left(\mathcal{F}^{\mathcal{L}}_{s,t}\right)_{s<t},\mathbb{P},\left(\theta_{h}\right)_{h})$ in the sense of 1.3.
Proof.
Let $\Omega$ and $\mathcal{F}^{\mathcal{L}}_{s,t}$ be as in 1.7. Define the measurable maps $\theta_{h}:\Omega\to\Omega$ by $(\theta_{h}\mathcal{L})_{s,t}(x,y):=\mathcal{L}_{s+h,t+h}(x,y)$. Properties $(1)$ and $(2)$ of 1.8 verify properties $(1)$ and $(2)$ of 1.3, respectively. Properties $(4)$ and $(5)$ of 1.3 are satisfied directly by the definition of the map $\theta_{h}$. Finally, we note that $\mathcal{F}^{\mathcal{L}}=\bigvee_{-\infty<s<t<\infty}\mathcal{F}^{\mathcal{L}}_{s,t}$ by definition. ∎
We can now state our main theorem.
Theorem 1.10.
The noise associated to the directed landscape is a black noise in the sense of 1.4.
There are a few equivalent conditions for demonstrating that a noise is a black noise (we use the one stated in 2.3). These conditions usually involve a variance bound which resembles a noise sensitivity condition with respect to all functions in $L^{2}(\Omega)$. Noise sensitivity, which was studied in the context of Boolean functions in foundational works such as [KKL88, BKS99], and then for its applications to physics and computer science in works such as [ST99, SS10, MOO10], is closely related to black noise. In fact, another way to understand the concept of black noise is to consider it as a system with respect to which all observables (functionals) are noise sensitive. It would be interesting to explore these noise sensitivity properties further and in a more precise fashion. Ganguly and Hammond [GH20], for instance, study the asymptotic noise sensitivity of Brownian last passage, using the notion of the overlap of perturbed geodesics to measure noise sensitivity.
While we do not study more specific noise sensitivity phenomena in this paper, we do obtain a few corollaries of 1.10. The next result, 1.11, demonstrates the independence of the directed landscape from any white noise defined on the same probability space. For the purpose of this text, a Gaussian white noise on an open subset $U\subset\mathbb{R}^{d}$ is a family of jointly Gaussian random variables $\{\xi(f)\}_{f\in L^{2}(U)}$ all defined on the same probability space, satisfying $\mathbb{E}[\xi(f)\xi(g)]=(f,g)_{L^{2}(U)}.$ We often write $\xi(f):=\int_{U}f(x)\xi(dx)$ or similar. A Gaussian space-time white noise is just a special case $U=\mathbb{R}^{2},$ where the first coordinate is interpreted as “time” and the second as “space.”
Theorem 1.11.
Let $(\Omega,(\mathcal{F}_{s,t})_{s<t},\mathbb{P},(\theta_{h})_{h})$ be a noise in the sense of 1.3. Let $\mathcal{L}$ and $\xi$ respectively be a copy of a directed landscape and a Gaussian space-time white noise, defined on this same probability space, which together generate the entire $\sigma$-algebra $\mathcal{F}:=\bigvee_{s<t}\mathcal{F}_{s,t}$, and which are both adapted to the filtration $(\mathcal{F}_{s,t})_{s<t}$, in the sense that
  • $\int_{\mathbb{R}^{2}}\phi(u,x)\xi(du,dx)$ is $\mathcal{F}_{s,t}$-measurable if $\phi$ is supported on $[s,t]\times\mathbb{R}.$
  • $\mathcal{L}_{a,b}(x,y)$ is $\mathcal{F}_{s,t}$ measurable for all $s\leq a<b\leq t$ and $x,y\in\mathbb{R}.$
Then $\mathcal{L}$ and $\xi$ are independent under $\mathbb{P}$.
The directed landscape is not unique in this respect: the proof will show that one could replace $\mathcal{L}$ by any space-time field that is known to be a black noise. This theorem may be surprising, since it illustrates a contrast with many prelimiting models in the KPZ class, in which case one often has $\mathcal{L}_{N}=F_{N}(\xi_{N})$ for some $(\mathcal{F}_{s,t})$-adapted path functionals $F_{N}$ of the discrete driving noise $\xi_{N}$. As we take the limit $N\to\infty$, 1.11 shows that this dependence no longer holds. In fact, the two fields $\mathcal{L}_{N}$ and $\xi_{N}$ become independent. The following corollary illustrates this phenomenon in the context of the KPZ equation.
Corollary 1.12 (The environment decouples from the height process when scaling to the landscape).
Let $\xi$ be a Gaussian space-time white noise on $\mathbb{R}_{+}\times\mathbb{R}$, and let $H^{\varepsilon}(t,x)$ denote the Hopf-Cole solution of the KPZ equation $\partial_{t}H^{\varepsilon}=\frac{1}{4}(\partial_{x}H^{\varepsilon})^{2}+\frac{\varepsilon}{4}\partial_{x}^{2}H^{\varepsilon}+\varepsilon^{1/2}\xi+\frac{1}{12}$, started from any deterministic sequence of upper semicontinuous initial data $H^{\varepsilon}(0,x)$ satisfying $H^{\varepsilon}(0,x)\leq C(1+|x|)$ for some $C>0$, and converging as $\varepsilon\to 0$ to some upper semicontinuous profile $H_{0}(x)$ in the sense of local hypograph convergence [MQR21, Section 3.1]. Consider any joint limit point $(\mathfrak{h},\xi)$ as $\varepsilon\to 0$ of the pair $(H^{\varepsilon},\xi)$ in the topology of $C((0,\infty)\times\mathbb{R})\times\mathcal{S}^{\prime}(\mathbb{R}^{2}).$ Then $\mathfrak{h}$ is distributed as the KPZ fixed point started from $H_{0}(x)$, independent of the space-time white noise $\xi.$
Before proving the 1.12, we note that joint limit points do exist because each coordinate is tight, which implies that the pair is tight. This theorem characterizes any joint limit point uniquely, and therefore it is a limit theorem for the pair $(H^{\varepsilon},\xi)$. The local hypograph convergence for the sequence of initial data is a weaker assumption than uniform convergence on compacts if all profiles are continuous.
Proof.
We define a four-parameter field $\mathscr{H}^{\varepsilon}_{s,t}(x,y)$ for each $\varepsilon>0$ such that $(t,y)\mapsto\mathscr{H}^{\varepsilon}_{s,t}(x,y)$ is the Hopf-Cole solution of the same equation solved by $H^{\varepsilon}$, with the same realization of the noise $\xi$, started at time $s$ from initial condition given by a narrow wedge at position $x$ (meaning that the exponential is started from a Dirac mass centered at location $x$). The main results of [QS20, Wu23] imply that this four-parameter field $\mathscr{H}^{\varepsilon}$ converges in law (in the topology of $C(\mathbb{R}^{4}_{\uparrow})$) to the directed landscape.
1.11 implies that any joint limit point of $(\mathscr{H}^{\varepsilon},\xi)$ must be of the form $(\mathcal{L},\xi)$ where $\mathcal{L}$ is a directed landscape independent of $\xi$. Indeed, one can define the $\sigma$-algebras $\mathcal{F}_{s,t}$ to be precisely the ones generated by random variables of the two forms given by the two bullet points in 1.11. The noise property needed to apply 1.11 holds in that limit simply because it holds trivially in the prelimit and any limit point will still respect this property.
Now we consider any joint limit point $(\mathfrak{h},\mathcal{L},\xi)$ of the triple $(H^{\varepsilon},\mathscr{H}^{\varepsilon},\xi)$. On one hand, we already know that $\mathcal{L}$ and $\xi$ must be independent under this limit point. On the other hand, the results of [NQR20, Wu23] imply that $\mathfrak{h}(t,y)=\sup_{x\in\mathbb{R}}H_{0}(x)+\mathcal{L}_{0,t}(x,y).$ Therefore $\mathfrak{h}$ is $\mathcal{L}$-measurable, which implies that it is independent of $\xi.$ Projecting onto the first and third marginals $\mathfrak{h}$ and $\xi$ gives the result. ∎
While 1.12 was formulated purely in terms of the KPZ equation, the result of 1.11 implies something much stronger: in any system for which there a height process which converges to the directed landscape, the underlying field of environment variables necessarily decouples from the height process under that limit. In the case of ASEP, the environment consists of the Poisson clocks that result in executed jumps for the system. In the case of last passage percolation or directed polymers, the underlying environment consists of the independent, identically distributed weights through which the last passage paths travel. 1.11 implies that for such systems, the environment and the height process asymptotically decouple as they respectively scale to the white noise and to the landscape [ACH24, DV21]. Such decoupling and sensitivity phenomena are not present in the intermediate-disorder or weak-asymmetry regime.
This phenomenon illustrates a contrast between the intermediate-disorder regime and the strong-disorder regime for $(1+1)$-dimensional systems in the KPZ universality class. For example, there are many “weak KPZ” scaling limit results where particle systems, SPDEs, or directed polymers are known to converge to the KPZ equation [BG97, AKQ14, DT16, DGP17, CST18, GJ14, HQ18, Yan23] as opposed to the KPZ fixed point. In these results, it is straightforward to show from the methods of those papers that the field of environment variables and the height process converge jointly in law to the driving noise of the limiting KPZ equation. This is always the case in the presence of intermediate disorder scaling or weakly asymmetric scaling of models in the KPZ university class. There are a few intuitive reasons why the same phenomenon does not happen when scaling to the directed landscape, i.e., in the strong disorder or fixed asymmetry regime. At the level of last passage percolation or directed polymers, the height process under strong disorder scaling only sees the largest variables in the environment field, whereas most of the smaller weights are ignored. The height process thus ignores “most” of the environment under the strong disorder scaling, with only a few special exceptions. In contrast, all of the noise variables contribute to some extent for the intermediate disorder scaling, as individual large contributions have a less dramatic effect on the overall height process.
We discuss another, perhaps more precise, conjecture about why the strong disorder scaling is so different from the intermediate disorder regime with regard to retaining the memory of the microscopic environment. As before, we work through the lens of polymers (for the purpose of this discussion, they may be discrete or continuous as the reader prefers). Polymers admit a chaos expansion in terms of multilinear polynomials of the underlying environment field. Distinct terms in this expansion are always orthogonal to one another in $L^{2}$ of the probability space associated to the underlying environment. Under intermediate disorder scaling, it is well understood from the work of Alberts, Khanin, and Quastel [AKQ14] that every individual term of the prelimiting chaos expansion contributes meaningfully to the scaling limit. In fact, they prove that the KPZ equation limit appears by proving term-by-term convergence of the polynomial chaos expansion of the polymer partition function to that of the exponential of the KPZ equation. In contrast, in the strong disorder regime, we conjecture that most of the $L^{2}$-mass of the terms of the chaos expansion of the height process escapes into the tails of the chaos expansion. And in fact, we conjecture that this “mass escape” phenomenon will not just occur for the height process itself, but also for every bounded measurable functional of the height process. The “mass escape” phenomenon is another way to formulate noise sensitivity of a sequence of observables [Gar11, Proposition 2.2].
As another corollary of 1.11, we show that the directed landscape is not an SPDE driven by space-time Gaussian white noise (in fact, any Gaussian noise that is white in time would not be sufficient to generate the landscape). To formalize the notion of being an SPDE, we note that any probabilistic strong solution of an SPDE defines an adapted path functional on the probability space of the white noise. We therefore show that no such adapted functional can exist.
Corollary 1.13 (The directed landscape is not an SPDE).
Let $\xi$ be a Gaussian space-time white noise on $\mathbb{R}^{2},$ defined on some probability space $(\Omega,\mathcal{F}^{\xi},\mathbb{P})$. For $s<t$ let $\mathcal{F}_{s,t}^{\xi}$ denote the $\sigma$-algebra generated by the random variables $\int_{\mathbb{R}^{2}}\phi(s,y)\xi(ds,dy)$ as $\phi$ ranges over all smooth functions of support contained in $[s,t]\times\mathbb{R}.$ Then there does not exist any random variable $\mathcal{L}$ taking values in $C(\mathbb{R}^{4}_{\uparrow})$, defined on the same probability space $\Omega$, satisfying the following properties:
  • $\mathcal{L}$ is distributed as the directed landscape.
  • $\mathcal{L}_{a,b}(x,y)$ is $\mathcal{F}^{\xi}_{s,t}$-measurable whenever $s\leq a<b\leq t$ and $x,y\in\mathbb{R}.$
The proof is immediate from 1.11, as $\mathcal{L}$ cannot be simultaneously independent of $\xi$ and adapted to its filtration: this would imply $\mathcal{L}$ is independent of itself and thus deterministic. We remark that this corollary is false without the adaptedness assumptions. Indeed a recent work of [DV24] shows that the landscape can indeed be written as some non-adapted functional of white noise, in fact through a measurable bijection.
Next, we discuss a few open questions that are related to the content of this paper. The two-dimensional critical stochastic heat flow is a continuum model recently constructed by [CSZ23a]. It is a flow of random measures on $\mathbb{R}^{2}$, written as $u_{s,t}(\vec{x},d\vec{y})$ indexed by $-\infty<s<t<\infty$ and $\vec{x},\vec{y}\in\mathbb{R}^{2}$. We can think of it as the random field obtained from the universal scaling limit of partition functions of intermediate-disorder directed polymers in a space-time random environment in spatial dimension $d=2$. This is another example of a random field which is a noise in the sense of 1.3, and it exhibits certain properties which strongly suggest that it will not be a white noise [CSZ23b].
Conjecture 1.14.
The two-dimensional critical stochastic heat flow constructed in [CSZ23a] is a black noise.
By the remark after 1.11, if this conjecture is true, it would be possible to prove decoupling theorems analogous to 1.12 for two-dimensional polymers. Our next conjecture is about the directed landscape being a noise in its space variables as well as its time variables. To formulate this precisely, we need to define the notion of being a two-dimensional noise. The following is adapted from the definition of a $d$-dimensional noise stated in Tsirelson [Tsi04], or in Ellis and Feldheim [EF16].
Definition 1.15.
A two-dimensional noise is a probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a collection of sub-$\sigma$-algebras $\mathcal{F}_{\vec{x},\vec{y}}\subset\mathcal{F}$ associated to all open two-dimensional rectangles in $\mathbb{R}^{2}$, and a collection of measurable maps $\theta_{\vec{h}}:\Omega\to\Omega$ indexed by $\vec{h}\in\mathbb{R}^{2}$, such that the following properties hold.
  1. (1)
    $\mathcal{F}_{\vec{x}_{1},\vec{y}_{1}}$ and $\mathcal{F}_{\vec{x}_{2},\vec{y}_{2}}$ are independent under $\mathbb{P}$ when $R_{1}:=(x_{1}^{(1)},x_{1}^{(2)})\times(y_{1}^{(1)},y_{1}^{(2)})$ and $R_{2}:=(x_{2}^{(1)},x_{2}^{(2)})\times(y_{2}^{(1)},y_{2}^{(2)})$ satisfy $R_{1}\cap R_{2}=\emptyset$.
  2. (2)
    $\mathcal{F}_{\vec{x}_{1},\vec{y}_{1}}\vee\mathcal{F}_{\vec{x}_{2},\vec{y}_{2}}=\mathcal{F}_{\vec{x}_{3},\vec{y}_{3}}$ whenever the rectangles $R_{1}$ and $R_{2}$ satisfy $R_{1}\cap R_{2}=\emptyset$ and $\overline{R_{1}\cup R_{2}}=\overline{R_{3}}.$
  3. (3)
    $\mathcal{F}$ is generated by the union of all of the $\mathcal{F}_{\vec{x},\vec{y}}$.
  4. (4)
    If $A\in\mathcal{F}_{\vec{x},\vec{y}}$, then $\theta_{\vec{h}}(A)\in\mathcal{F}_{\vec{x}+(h_{1},h_{1}),\vec{y}+(h_{2},h_{2})}$ and $\mathbb{P}(\theta_{\vec{h}}(A))=\mathbb{P}(A)$.
  5. (5)
    $\theta_{\vec{h}_{1}}\theta_{\vec{h}_{2}}=\theta_{\vec{h}_{1}+\vec{h}_{2}}$ for all $\vec{h}_{1},\vec{h}_{2}\in\mathbb{R}^{2}$, and $\theta_{\vec{0}}=\text{Id}.$
We say this two-dimensional noise is a black noise when the only linear random variable is 00, i.e., any $F\in L^{2}(\Omega)$ satisfying $\mathbb{E}[F|\mathcal{F}_{R_{1}\cup R_{2}}]=\mathbb{E}[F|\mathcal{F}_{R_{1}}]+\mathbb{E}[F|\mathcal{F}_{R_{2}}]$ for all rectangles $R_{1},R_{2}$ as in Item (2), implies that $F=0$.
[…]
Conjecture 1.16.
The directed landscape is a two-dimensional black noise.
The difficulty of the two-dimensional problem is showing that the noise property holds in the first place, in particular Item (2) in 1.15. If true, this would address a problem posed by Ellis and Feldheim [EF16]. Their work demonstrates that the Brownian web is an example of a two-dimensional black noise, and they ask for further relevant examples of two-dimensional black noise arising from simple probabilistic systems.
There are several key properties of the directed landscape that are crucial to the proof that it is a black noise, and which distinguish it from many prelimiting models in the KPZ universality class that are inherently white noise.
  1. (1)
    Scale invariance: Obtaining a rescaled version of the same space-time process at every temporal scale is a crucial ingredient to many parts of the proof. For example, in 3.3 we get a spatial mixing rate for the directed landscape of order $e^{-dk^{3}}$ on a time interval $[0,1]$. This mixing rate immediately translates to a mixing rate of $e^{-d\epsilon^{-2}k^{3}}$ for on a smaller time interval $[0,\epsilon]$.
  2. (2)
    Zero Temperature: Many of our estimates would be weakened or fail for positive temperature analogues like the KPZ equation, as the maximum in the definition of the directed landscape would get replaced by a convolution operation which is not as local in some sense.
  3. (3)
    Local Brownianity: Even for many zero-temperature conjectural prelimits (such as inviscid stochastic Burgers equations driven by white noise) the local Brownianity of the spatial process fails. Improved spatial regularity of the solution actually hurts many of the estimates one needs to prove the black noise property, as the better regularity “widens” the set of $x$-values where the spatial height process is close to the argmax. Controlling the size of the latter set of $x$-values is really the heart of the proof, as quantified in 6.4.
We also remark that the black noise property seems intimately related to the degeneracy and fractal structure of the landscape, for instance the tendency of geodesics to overlap. The papers [CHHM22, GZ22, Dau24, DG21] study this fractal structure in some depth. It remains an open problem to precisely and quantitatively relate the black noise property of the landscape to its fractal structure. For instance, the optimal exponents in the variance bounds needed to prove the black noise property should be related to the Hausdorff dimension of the endpoints of exceptional geodesics.
1.3. Notation
Throughout the paper, we use the font $\mathfrak{B},\mathfrak{C}$ as a way to denote random constants. The letters $C,d$ are used, respectively, to denote deterministic constants that may get larger or smaller from line to line when performing an estimate. Typically, such $C,d$ will be (respectively) a multiplying constant and an exponential decay rate of some quantity. For a metric space $X$ we will let $C(X)$ denote the space of continuous functions $X\to\mathbb{R}$, equipped with the topology of uniform convergence on compacts. We often deal with random variables defined on the underlying probability space $C(\mathbb{R})$ or $C(\mathbb{R}^{2})$ or $C(\mathbb{R}^{4}_{\uparrow})$, and these random variables should be understood as being real-valued Borel-measurable functions with respect to the topology of uniform convergence on compact sets (which is always separable and metrizable, hence Polish).
1.4. Structure of the Paper
In Section 2, we reduce the problem of demonstrating that the directed landscape is a black noise to demonstrating a specific variance estimate about the Airy sheet (2.6). Section 3 through Section 6 are dedicated to obtaining this estimate, which relies on a mixing property of the directed landscape (Section 3), as well as estimates on argmaxes of Airy sheets (Section 4) and estimates on Bessel processes (Section 5). Then in Section 6, we conclude that the directed landscape is a black noise. Finally in Section 7, we prove 1.11 from 1.10.
1.5. Acknowledgements
SP thanks Duncan Dauvergne, Jeremy Quastel, and Bálint Virág for discussions and context about this problem during a visit to Toronto. SP and ZH thank Ivan Corwin and Yu Gu for comments on an earlier draft of the paper. ZH was supported by the Fernholz Foundation’s Minerva Fellowship Program and Ivan Corwin’s grant, NSF DMS-1811143.
2. Reducing the black noise problem to an estimate on the Airy sheet
In this section, we reduce the problem of showing that the directed landscape is a black noise to a specific variance bound for the Airy sheet.
Lemma 2.1.
[…]
Proof.
We define $\mathcal{G}_{m}:=\sigma(X_{1},...,X_{m})$ and $F_{m}:=\mathbb{E}[F|\mathcal{G}_{m}].$ Using approximation by $\mathcal{G}_{m}$-measurable simple functions, we see that $F_{m}$ are of the form $\Psi_{m}(X_{1},...,X_{m})$ for some Borel measurable $\Psi_{m}:\mathcal{X}^{m}\to\mathbb{R}$. If needed, we replace $\Psi_{m}$ by $(-\|F\|_{L^{\infty}(\Omega)})\vee\Psi_{m}\wedge\|F\|_{L^{\infty}(\Omega)}$, and this still satisfies the same relation $F_{m}=\Psi_{m}(X_{1},...,X_{m})$ on a set of measure $1$. By Doob’s martingale convergence theorem, $F_{m}\to F$ almost surely and in every $L^{p}(\Omega)$. ∎
In his survey, Tsirelson [Tsi04] gives a few examples in which it is possible to prove something is a black noise, and provides a necessary and sufficient condition for a given noise to be black. We quote the following result from that survey, see [Tsi04, Proposition 7a(3) Item (b)].
Proposition 2.2.
Let $(\Omega,(\mathcal{F}_{s,t})_{s<t},\mathbb{P},\theta_{h})$ be a noise in the sense of 1.3. The space $M$ of $\mathcal{F}_{0,1}$-measurable linear random variables (1.4) is a closed linear subspace of $L^{2}(\Omega)$, and moreover the orthogonal projection of any $F\in L^{2}(\Omega)$ onto $M$ is given by
$\lim_{n\to\infty}\sum_{k=1}^{2^{n}}\big{(}\mathbb{E}\left[F\big{|}\mathcal{F}_{(k-1)2^{-n},k2^{-n}}\right]-\mathbb{E}[F]\big{)},$
where the limit is understood in $L^{2}(\Omega).$
Corollary 2.3.
A noise $(\Omega,(\mathcal{F}_{s,t})_{s<t},\mathbb{P},\theta_{h})$ is a black noise if and only if the following condition holds on a dense linear subspace $D\subset L^{2}(\Omega)$:
$\sum_{k=1}^{2^{n}}\big{(}\mathbb{E}\left[F\big{|}\mathcal{F}_{(k-1)2^{-n},k2^{-n}}\right]-\mathbb{E}[F]\big{)}\stackrel{{\scriptstyle L^{2}(\Omega)}}{{\to}}0,\;\;\;\;\;\;\;\;\mathrm{for\;\;all\;\;\;}F\in D,$
or equivalently,
$\lim_{n\to\infty}\sum_{k=1}^{2^{n}}\mathrm{Var}\left(\mathbb{E}\left[F\big{|}\mathcal{F}_{(k-1)2^{-n},k2^{-n}}\right]\right)=0,\;\;\;\;\;\;\;\;\mathrm{for\;\;all\;\;\;}F\in D.$ (2.1)
We recall the definition of the directed landscape and the associated noise (1.8) and use $\Omega$ to denote the underlying probability space of $\mathcal{L}$. There is a particular class of functions that are most useful to us in the context of the directed landscape. We define $D\subset L^{2}(\Omega)$ to be those functionals of the landscape that are finite linear combinations of “polynomial functions” of the form
$F=\prod_{j=1}^{m}\prod_{i=1}^{n_{j}}\mathcal{L}_{s_{j},t_{j}}(x_{ij},y_{ij}),$ (2.2)
where $(s_{j},t_{j})$ are disjoint intervals of $\mathbb{R}$ whose endpoints $s_{j},t_{j}$ are both dyadic numbers (of the form $k2^{-N}$ with $k\in\mathbb{Z}$ and $N\in\mathbb{N}),$ furthermore $x_{ij},y_{ij}\in\mathbb{R}$ are all distinct points in $\mathbb{R}$, and $m,n_{j}\in\mathbb{N}$.
Lemma 2.4.
The set $D$ is a dense linear subspace of $L^{2}(\Omega)$ with $\Omega$ as in 1.7.
Proof.
We recall the following tensorization property of $L^{2}$ spaces: the space $L^{2}(\gamma_{1})\otimes\cdots\otimes L^{2}(\gamma_{m})$ is isometrically isomorphic as a Hilbert space to $L^{2}(\gamma_{1}\otimes\cdots\otimes\gamma_{m})$ for any Borel probability measures $\gamma_{i}$ on a Polish space, under the canonical isomorphism $(F_{1}\otimes\cdots\otimes F_{m})(x_{1},...,x_{m})\cong F_{1}(x_{1})\cdots F_{m}(x_{m})$. Therefore, in order to show that functions of the form (2.2) are dense in $L^{2}(\Omega)$, it suffices to show that any function of the form $\prod_{j=1}^{m}\varphi_{j}(\mathcal{L}_{s_{j},t_{j}}(x_{1},y_{1}),...,\mathcal{L}_{s_{j},t_{j}}(x_{n},y_{n}))$ is in the $L^{2}$-closure of $D$, where $m,n\in\mathbb{N}$ and $\varphi_{j}$ are bounded measurable functions on $\mathbb{R}^{n}$. Indeed, functionals of the latter form are dense in $L^{2}(\Omega)$ by the following argument: by 2.1 we can take $n\uparrow\infty$ to approximate any functional of the form $\prod_{j=1}^{m}\Phi_{j}(\mathcal{L}_{s_{j},t_{j}})$ where $m\in\mathbb{N}$, and $\Phi_{j}:C(\mathbb{R}^{2})\to\mathbb{R}$ are bounded measurable, and $\mathcal{L}_{s,t}:\mathbb{R}^{2}\to\mathbb{R}$ denotes the random function given by $(x,y)\mapsto\mathcal{L}_{s,t}(x,y)$. Then in turn by taking linear combinations and using the tensorization property of $L^{2}$ we can approximate any functional of the form $\Psi(\mathcal{L}_{s_{1},t_{1}},...,\mathcal{L}_{s_{m},t_{m}})$ where $m\in\mathbb{N}$ and $\Psi:C(\mathbb{R}^{2})^{m}\to\mathbb{R}$ is bounded measurable, and finally by 2.1 we can take $m\uparrow\infty$ and the latter types of functionals can then approximate any general functional $F(\mathcal{L}).$
We therefore focus on showing that any function of the form $\prod_{j=1}^{m}\varphi_{j}(\mathcal{L}_{s_{j},t_{j}}(x_{1},y_{1}),...,\mathcal{L}_{s_{j},t_{j}}(x_{n},y_{n}))$ is in the $L^{2}(\Omega)$-closure of $D$, where $m,n\in\mathbb{N}$ and $\varphi_{j}$ are bounded measurable functions on $\mathbb{R}^{n}$. By independence of the different terms in the product, it suffices to prove that, for $m=1$, such functions may be approximated in $L^{2}(\Omega)$ by functions of the form (2.2) with $m=1$. In turn, it suffices to show that polynomial functions on $\mathbb{R}^{n}$ are dense in $L^{2}(\mu)$ where $\mu$ is the probability measure on $\mathbb{R}^{n}$ given by the joint law of $(\mathcal{L}_{s,t}(x_{1},y_{1}),...,\mathcal{L}_{s,t}(x_{n},y_{n}))$ for some fixed interval $(s,t)$ and fixed $\vec{x},\vec{y}\in\mathbb{R}^{n}$. Since the marginals here are deterministically shifted Tracy-Widom GUE laws which have a super-exponential tail decay, we conclude that $\mu$ must have a finite moment generating function everywhere. Therefore, if $f\in L^{2}(\mu)$ was orthogonal to all polynomials, then by taking the Taylor expansion of the exponential and interchanging the infinite sum with the integral over $\mathbb{R}^{n}$, we would have that $\int_{\mathbb{R}^{n}}f(x)e^{i\sum_{1}^{n}\lambda_{j}x_{j}}\mu(dx)=0$ for all $\lambda=(\lambda_{1},...,\lambda_{n})\in\mathbb{R}^{n}$. This interchange can be justified by Fubini, using the super-exponential tail decay on $\mu$. By inverting the Fourier transform (which is a linear isomorphism on $\mathcal{S}^{\prime}(\mathbb{R}^{n}))$, this implies that the measure $f\;d\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$ must be the zero measure, i.e., $f=0$ $\mu$-almost everywhere. ∎
Due to the independence of the $\sigma$-algebras associated to disjoint intervals (property $(1)$ in 1.8), we can ignore the product over $j$ in (2.2), and thus without loss of generality, we may assume that $m=1$ in (2.2). We label the single interval by $(s_{1},t_{1})=(s,t)$. By shifting and rescaling the landscape, we can assume without loss of generality that $(s,t)=(0,1).$ Therefore, we have reduced the black noise problem to demonstrating that
$\lim_{n\to\infty}\sum_{k=1}^{2^{n}}\mathrm{Var}\left(\mathbb{E}\left[\prod_{i=1}^{n}\mathcal{L}_{0,1}(x_{j},y_{j})\bigg{|}\mathcal{F}^{\mathcal{L}}_{(k-1)2^{-n},k2^{-n}}\right]\right)=0,$ for $\text{for all }\vec{x},\vec{y}\in\mathbb{R}^{n},\text{ and }n\in\mathbb{N}.$ (2.3)
The entire problem is now formulated in terms of the Airy sheet $\mathcal{S}:=\mathcal{L}_{0,1}$. In fact, we can simplify things further.
Proposition 2.5.
In order to show that the directed landscape is a black noise, it suffices to show that for any $\eta\in(0,1/2)$ and $n\in\mathbb{N}$ and $\vec{x},\vec{y}\in\mathbb{R}^{n}$,
$\lim_{n\to\infty}\sum_{\eta 2^{n}\leq k\leq(1-\eta)2^{n}}\mathrm{Var}\left(\mathbb{E}\left[\prod_{i=1}^{n}\mathcal{L}_{0,1}(x_{j},y_{j})\bigg{|}\mathcal{F}^{\mathcal{L}}_{(k-1)2^{-n},k2^{-n}}\right]\right)=0,$ (2.4)
Proof.
Let $F:=\prod_{i=1}^{n}\mathcal{L}_{0,1}(x_{j},y_{j}),$ and let $G$ denote the linear random variable obtained by applying the orthogonal projection of $F$ onto the closed linear subspace of $L^{2}(\Omega)$ consisting of linear $\mathcal{F}^{\mathcal{L}}_{0,1}$ measurable random variables.
By applying 2.2, we see from (2.4) that $G$ must be measurable with respect to $\mathcal{F}^{\mathcal{L}}_{0,\eta}\vee\mathcal{F}^{\mathcal{L}}_{1-\eta,1}$, in other words, all dependency on $\mathcal{F}^{\mathcal{L}}_{\eta,1-\eta}$ vanishes. This is true for all $\eta\in(0,1/2)$. Therefore, $G$ is measurable with respect to $\mathcal{H}:=\bigcap_{\eta\in(0,1/2)}\mathcal{F}^{\mathcal{L}}_{0,\eta}\vee\mathcal{F}^{\mathcal{L}}_{1-\eta,1}.$ We claim that $\mathcal{H}$ is a trivial $\sigma$-algebra, only containing sets of measure 00 or $1$. It is clear that $\mathcal{H}$ is contained in $\mathcal{F}^{\mathcal{L}}_{0,1}.$ However, $\mathcal{H}$ is also independent of the $\sigma$-algebra generated by $\bigcup_{\eta\in(0,1/2)}\mathcal{F}^{\mathcal{L}}_{\eta,1-\eta}.$ We claim that the $\sigma$-algebra generated by $\bigcup_{\eta\in(0,1/2)}\mathcal{F}^{\mathcal{L}}_{\eta,1-\eta}$ is all of $\mathcal{F}^{\mathcal{L}}_{0,1}.$ To see this, note by the continuity of the directed landscape in all parameters that e.g. the random variable $\mathcal{L}_{0,1}(x,y)$ can be written as the almost sure limit as $\eta\to 0$ of $\mathcal{L}_{\eta,1-\eta}(x,y).$ Thus $\mathcal{H}\subset\mathcal{F}^{\mathcal{L}}_{0,1}$ is independent of $\mathcal{F}^{\mathcal{L}}_{0,1}$, proving that it is a trivial $\sigma$-algebra.
Since the random variable $G$ is $\mathcal{H}$-measurable, we conclude that it is almost surely constant. But $G$ is an $L^{2}$-limit of the sequence of random variables as in 2.2, and therefore (2.3) holds. ∎
The following corollary is an immediate consequence.
Corollary 2.6.
Fix some $\varrho>0$. In order to show that the directed landscape is a black noise, it suffices to show that for any $\eta\in(0,1/2)$, $n\in\mathbb{N}$ and $\vec{x},\vec{y}\in\mathbb{R}^{n}$, there exists some $C=C(n,\eta,\vec{x},\vec{y})>0$ such that for all $a,b\in[\eta,1-\eta]$,
$\mathrm{Var}\left(\mathbb{E}\left[\prod_{i=1}^{n}\mathcal{L}_{0,1}(x_{j},y_{j})\bigg{|}\mathcal{F}^{\mathcal{L}}_{a,b}\right]\right)\leq C(b-a)^{1+\varrho},$ (2.5)
The rest of the paper will focus on proving (2.5), for a value of $\varrho$ that is slightly smaller than $\frac{1}{15}$. This bound may not be sharp.
3. A strong mixing property for the landscape under spatial shifts
In this section, we prove a strong mixing property for the directed landscape under spatial shifts (3.5) that is crucial to the arguments in Section 4. First, we state a theorem from [DOV22] that is used in the proof.
Theorem 3.1 (Theorem 1.7 [DOV22]).
Fix $u=(x,t;y,s)\in\mathbb{R}^{4}_{\uparrow}$. There is almost surely a unique directed geodesic $\Pi_{u}$ from $(x,s)$ to $(y,t)$, i.e., a path $\pi:[s,t]\to\mathbb{R}$ such that $\pi(s)=x,\pi(t)=y$ and for all $k\in\mathbb{N}$ and all $s=s_{0}<...<s_{k}=t$,
$\mathcal{L}_{s,t}(x,y)=\sum_{j=1}^{k}\mathcal{L}_{t_{k-1},t_{k}}(\pi(t_{k-1}),\pi(t_{k})).$
Its distribution only depends on $u$ through scaling: as random continuous functions from $[0,1]\to\mathbb{R}$,
[…]
Moreover, for $u=(0,0;0,1)$, there is a random constant $\mathfrak{C}$ such that, for all $s>0$ with $t,t+s\in[0,1]$,
$\big{|}\Pi_{u}(t+s)-\Pi_{u}(t)\big{|}\leq\mathfrak{C}s^{\frac{2}{3}}\log^{\frac{1}{3}}{\left(\frac{2}{s}\right)}.$
The random constant satisfies $\mathbb{E}\big{[}a^{\mathfrak{C}^{3}}\big{]}<\infty$ for some $a>1$.
With 3.1 in hand, we prove that there exists a probability space on which we can couple three copies of the directed landscape in such a way that two of them are independent, and the third agrees with the other two sufficiently far from the origin.
Lemma 3.2.
There exist three copies of the directed landscape $\mathcal{L}^{i}=(\mathcal{L}^{i}_{s,t})_{0\leq s<t\leq 1}$ with $i=0,1,2$, all coupled onto the same probability space $(\Omega,\mathcal{F},\mathbb{P}_{\mathrm{triple}})$ so that:
  1. (1)
    $\mathcal{L}^{1},\mathcal{L}^{2}$ are independent under $\mathbb{P}_{\mathrm{triple}}$.
  2. (2)
    There exists a positive random variable $X:\Omega\to(0,\infty)$ such that $\mathcal{L}_{s,t}^{0}(x,y)=\mathcal{L}_{s,t}^{1}(x,y)$ for all $x,y>X$ and $0\leq s<t\leq 1$, and furthermore $\mathcal{L}_{s,t}^{0}(x,y)=\mathcal{L}_{s,t}^{2}(x,y)$ for all $x,y<-X$ and $0\leq s<t\leq 1$.
  3. (3)
    $\mathbb{P}_{\mathrm{triple}}(X>u)\leq Ce^{-du^{3}}$ where $C,d>0$ do not depend on $u>0$.
This lemma is inspired by [Dau24, Proposition 2.6], which proves a coupling result for copies of the directed landscape via approximation by a last passage model.
Proof of 3.2.
We consider two fields of independent, identically distributed exponential random variables $\omega^{i}=\{\omega^{i}_{t,x}\}_{t\geq 0,x\in\mathbb{Z}}$ for $i=1,2$, independent of one another. We define a third field of independent, identically distributed variables $\omega^{0}_{t,x}$ by $\omega^{0}_{t,x}:=\omega^{1}_{t,x}$ if $x\geq 0$ and $\omega^{0}_{t,x}:=\omega^{2}_{t,x}$ if $x<0$.
For $i=0,1,2$ and $x,y\in\mathbb{Z}$ and $0\leq m<n\in\mathbb{Z}$, we consider the fields of last passage times
$L_{m,n}^{i}(x,y):=\max\{W^{i}(\pi):\pi:\{m,...,n\}\to\mathbb{Z};\pi(m)=x,\pi(n)=y\}$
where $\pi$ denotes a nearest neighbor path $\pi:\{m,...,n\}\to\mathbb{Z}$ and its length is defined by
$W^{i}(\pi):=\sum_{i=m}^{n}\omega^{i}_{i,\pi(i)}.$
For $x,y\in\varepsilon^{2/3}\mathbb{Z}$ and $s,t\in\varepsilon\mathbb{Z}$, we then define a rescaled last passage time
$\mathcal{L}^{i,\varepsilon}_{s,t}(x,y):=\varepsilon^{1/3}\big{(}L_{\varepsilon^{-1}s,\varepsilon^{-1}t}^{i}(\varepsilon^{-2/3}x,\varepsilon^{-2/3}y)-\varepsilon^{-1}(t-s)\big{)}.$
[DV21, Theorem 1.20] shows that the marginal laws of each $\mathcal{L}^{i,\varepsilon}_{s,t}(x,y)$ individually converge in law as $\varepsilon\to 0$ to the directed landscape in the uniform-on-compact topology. Here, the compact set can be taken to be any compact subset of $\{(s,t,x,y):0\leq s<t\leq 1,(x,y)\in\mathbb{R}^{2}\},$ see 1.7.
Furthermore, if $\pi_{\mathrm{opt}}^{i,m,n,x,y}$ denotes the maximizing path, i.e., satisfying $L^{i}_{m,n}(x,y)=W^{i}(\pi_{\mathrm{opt}}^{i,m,n,x,y})$, then [DV21, Theorem 8.7] implies that the paths
$\Pi^{i,s,t,x,y;\varepsilon}(r):=\varepsilon^{2/3}\pi_{\mathrm{opt}}^{i,\varepsilon^{-1}s,\varepsilon^{-1}t,\varepsilon^{-2/3}x,\varepsilon^{-2/3}y}(\varepsilon^{-1}r),$
converge in law as $\varepsilon\to 0$, jointly in all variables $(s,t,x,y,r)$, with respect to the uniform-on-compact topology on $\{(s,t,x,y,r):0\leq s<t\leq 1,s\leq r\leq t,(x,y)\in\mathbb{Z}^{2}\}$, to geodesic paths $\Pi^{i,s,t,x,y}(r)$ of the corresponding directed landscapes $\mathcal{L}^{i}$, for $i=0,1,2$. This convergence occurs jointly with the height profiles $\mathcal{L}^{i,\varepsilon}_{s,t}(x,y)$.
Now let $(\mathcal{L}^{0},\mathcal{L}^{1},\mathcal{L}^{2})$ be any joint limit point of $(\mathcal{L}^{0,\varepsilon},\mathcal{L}^{1,\varepsilon},\mathcal{L}^{2,\varepsilon})$ as $\varepsilon\to 0.$ We define the random variable
$X:=\inf\{x\in\mathbb{N}:\inf_{r\in[0,1]}\Pi^{0,0,1,x,x}(r)>0,\sup_{r\in[0,1]}\Pi^{0,0,1,-x,-x}(r)<0\},$
which tracks the smallest positive integer value of $x$ for which both of the directed geodesic paths from $x\to x$ and from $-x\to-x$ do not cross 00 in $\mathcal{L}^{0}$. We verify that $\mathcal{L}^{1},\mathcal{L}^{2}$ are independent and that $\mathcal{L}_{s,t}^{0}(x,y)=\mathcal{L}_{s,t}^{1}(x,y)$ for all $x,y>X$ and $0\leq s<t\leq 1$, by noting that the analogous facts hold trivially in the prelimit. Furthermore, $\mathcal{L}_{s,t}^{0}(x,y)=\mathcal{L}_{s,t}^{2}(x,y)$ for all $x,y<-X$ and $0\leq s<t\leq 1$, again because the analogous fact holds in the prelimit. This is due to the non-crossing property of geodesics.
It remains to verify the tail bound in Item (3) of the lemma statement. The moment bound of 3.1 implies that for any fixed $x\in\mathbb{R}$ the probability of any geodesic path from $x\to x$ varying more than a distance $u>0$ from its starting point decays like $Ce^{-du^{3}}$ for some $C,d>0$. This directly implies Item $(3)$ through a union bound. ∎
We use 3.2 to prove a strong mixing property for the landscape under diagonal shifts.
Proposition 3.3.
The directed landscape $\mathcal{L}=(\mathcal{L}_{s,t})_{0\leq s<t\leq 1}$ is $\alpha$-mixing under spatial shifts, with mixing rate $\alpha(k)\leq Ce^{-dk^{3}}$ for some universal constants $C,d>0$. More precisely, we have the following stronger estimate. For $-\infty\leq a<b\leq\infty$ let $\mathcal{G}_{a,b}$ denote the $\sigma$-algebra generated by the random variables $\mathcal{L}_{s,t}(x,y)$ where $0\leq s<t\leq 1$ and $x,y\in(a,b)$. Then
$|\operatorname{Cov}(F,G)|\leq Ce^{-dk^{3}}\mathbb{E}[F^{4}]^{1/4}\mathbb{E}[G^{4}]^{1/4},$
where $C,d>0$ are independent of $k$ and $F,G\in L^{4}(\Omega)$ that are respectively measurable with respect to $\mathcal{G}_{-\infty,0}$ and $\mathcal{G}_{k,\infty}.$
The $\alpha$-mixing property is a very strong form of mixing that is uniform over all observables (see the survey by Bradley [Bra05] for further explanation). Roughly speaking, the result says that under translation of the landscaped by $k$ spatial units, the decay of correlation is upper bounded by $e^{-dk^{3}}$, where $d>0$ is bounded away from zero uniformly over all observables and $k>0$. Moreover, the remarkably fast cubic-exponential rate of decay means that we can treat the field as virtually independent on well-separated scales.
Proof.
By translation invariance, it suffices to prove the claim with the assumption “$F$ measurable with respect to $\mathcal{G}_{-\infty,0}$” replaced by the statement “$F$ measurable with respect to $\mathcal{G}_{-\infty,-k}$” (this replacement is just for notational convenience). Therefore, we will consider $F,G$ that are respectively measurable with respect to $\mathcal{G}_{-\infty,-k}$ and $\mathcal{G}_{k,\infty}.$
We assume that $\mathcal{L}^{0},\mathcal{L}^{1},\mathcal{L}^{2}$ are the same couplings and $X$ is the same random variable as in 3.2. We will use $\mathbb{E}_{\mathrm{triple}}[\cdot]$ to denote the expectation with respect to the measure on the coupled space from 3.2. Consider the measurable functions $F,G$ as above and write $F_{i}:=F(\mathcal{L}^{i}),G_{i}:=G(\mathcal{L}^{i})$ for $i=0,1,2$. Then by independence, $\mathbb{E}_{\text{triple}}[F_{2}G_{1}]=\mathbb{E}_{\text{triple}}[F_{2}]\mathbb{E}_{\text{triple}}[G_{1}],$ thus
$|\operatorname{Cov}(F,G)|$ $=|\mathbb{E}_{\text{triple}}[F_{0}G_{0}-F_{2}G_{1}]|$
$\leq|\mathbb{E}_{\text{triple}}[(F_{0}-F_{2}+F_{2})(G_{0}-G_{1}+G_{1})-F_{2}G_{1}]|$
$=|\mathbb{E}_{\text{triple}}[(F_{0}-F_{2})(G_{0}-G_{1})]+\mathbb{E}_{\text{triple}}[(F_{0}-F_{2})G_{1}]+\mathbb{E}_{\text{triple}}[F_{2}(G_{0}-G_{1})]|$
$\leq\mathbb{E}_{\text{triple}}[(F_{0}-F_{2})^{2}]^{1/2}\mathbb{E}_{\text{triple}}[(G_{0}-G_{1})^{2}]^{1/2}+\mathbb{E}_{\text{triple}}[(F_{0}-F_{2})^{2}]^{1/2}\mathbb{E}_{\text{triple}}[G_{1}^{2}]^{1/2}$
$\;\;\;\;\;\;+\mathbb{E}_{\text{triple}}[F_{2}^{2}]^{1/2}\mathbb{E}_{\text{triple}}[(G_{0}-G_{1})^{2}]^{1/2}.$
Now recall that $F,G$ are, respectively, measurable with respect to $\mathcal{G}_{-\infty,-k}$ and $\mathcal{G}_{k,\infty}$. Therefore, $F_{0}=F_{2}$ and $G_{0}=G_{1}$ on the event that $X\leq k$. By 3.2 we thus have that
$\mathbb{E}_{\text{triple}}[(F_{0}-F_{2})^{2}]^{1/2}$ $=\mathbb{E}_{\text{triple}}[(F_{0}-F_{2})^{2}1_{\{X>k\}}]^{1/2}$
$\leq 2\mathbb{E}[F^{4}]^{1/4}\mathbb{P}_{\text{triple}}(X>k)^{1/2}\leq 2\mathbb{E}[F^{4}]^{1/4}Ce^{-\frac{1}{2}dk^{3}}.$
We used Cauchy-Schwarz and $(F_{0}-F_{2})^{2}\leq 2(F_{0}^{2}+F_{2}^{2})$ in the first inequality. Likewise, we may prove that $\mathbb{E}_{\text{triple}}[(G_{0}-G_{1})^{2}]^{1/2}\leq 2\mathbb{E}[G^{4}]^{1/4}Ce^{-\frac{1}{2}dk^{3}}$. On the other hand, it is clear from Jensen that $\mathbb{E}_{\text{triple}}[F_{2}^{2}]^{1/2}\leq\mathbb{E}_{\text{triple}}[F_{2}^{4}]^{1/4}=\mathbb{E}[F^{4}]^{1/4}$ and likewise $\mathbb{E}_{\text{triple}}[G_{1}^{2}]^{1/2}\leq\mathbb{E}_{\text{triple}}[G_{1}^{4}]^{1/4}=\mathbb{E}[G^{4}]^{1/4}$. ∎
The choice of $L^{4}$ is not optimal in this proof, and applying Hölder’s inequality rather than Cauchy-Schwarz could improve the assumptions to $F,G\in L^{p}(\Omega)$ with a bound $\mathbb{E}[|F|^{p}]^{1/p}\mathbb{E}[|G|^{p}]^{1/p}$ for any $p>2$. It is natural to ask what the strongest or optimal form of mixing is for spatial shifts in the landscape, such as $\beta$-mixing or $\phi$-mixing [Bra05], but we will not pursue this question.
Although 3.2 and 3.3 provide a flavor of how to prove mixing estimates, we will actually need a stronger version of these results.
Theorem 3.4.
Fix $T>0$, and consider any compact disjoint intervals $I_{1},J_{1},...,I_{N},J_{N}\subset\mathbb{R}$. Define the $\sigma$-fields
$\mathcal{F}_{I}:=\sigma(\{\mathcal{L}_{s,t}(x,y):0\leq s<t\leq T,x\text{ and }y\text{ are in the same interval }I_{r}\text{ for some }r\leq N\}$
$\mathcal{F}_{J}:=\sigma(\{\mathcal{L}_{s,t}(x,y):0\leq s<t\leq T,x\text{ and }y\text{ are in the same interval }J_{r}\text{ for some }r\leq N\}.$
Define the positive real number
$D:=\inf\bigg{\{}|x-y|:x\in\bigcup_{1}^{N}I_{r},\;\;y\in\bigcup_{1}^{N}J_{r}\bigg{\}}.$
Then we have the bound
$|\operatorname{Cov}(F,G)|\leq CNe^{-dT^{-2}D^{3}}\mathbb{E}[F^{4}]^{1/4}\mathbb{E}[G^{4}]^{1/4}.$
Here $C,d>0$ are universal constants that are independent of $N,T,\{I_{r},J_{r}\}_{r=1}^{N},D$ and $F,G\in L^{4}(\Omega)$ that are respectively measurable with respect to $\mathcal{F}_{I}$ and $\mathcal{F}_{J}.$
We remark that not all $I_{r}$ need to be strictly to the left of all $J_{r}$ or vice versa: they may very well be mixed amongst each other in a nontrivial way.
Proof.
Without loss of generality, as in Section 2, we may assume that $T=1$. By scale invariance of the directed landscape 1.8 (4), all of the intervals could be replaced by their $T^{-2/3}$-rescalings. This changes the value of $D$ to $T^{-2/3}D$, which respects the estimate. In the case that $T=1$, the proof uses an explicit coupling construction that is very similar to the proofs of 3.2 and 3.3. We now sketch the details of this coupling construction.
Without loss of generality, we assume that the intervals are ordered $I_{1},J_{1},I_{2},J_{2},...,I_{N},J_{N}$ so that the supremum of $I_{1}$ is less than or equal to the infimum of $J_{1}$, the supremum of $J_{1}$ is less than the infimum of $I_{2}$, and so on. This does not lose any generality because we can always combine consecutive instances of the $I_{r}$ or $J_{r}$ without changing the value of $D$.
We relabel $I_{1},J_{1},I_{2},J_{2},...,I_{N},J_{N}$ as $U_{1},...,U_{2N}$ and define $a_{r}$ to be the midpoint of the segment in the complement of $\bigcup_{i\leq 2N}U_{i}$ that is between $U_{r}$ and $U_{r+1}$. We define a fattening of $I_{r}$ and $J_{r},$ by setting $\bar{I}_{r}:=(a_{2r-1},a_{2r})$ and $\bar{J}_{r}:=(a_{2r},a_{2r+1})$ for $r=1,...,N$. In this notation, we say that $a_{2N+1}:=\infty$. We also define $b_{r},c_{r}$ to be those points in the segment in the complement of $\bigcup_{i\leq 2N}U_{i}$ that is between $U_{r}$ and $U_{r+1}$ that are (respectively) exactly one-third and two-thirds of the way across this segment.
If $L^{1},L^{2}\in C(\mathbb{R}_{\uparrow}^{4})$ and $I\subset\mathbb{R}$, then we will say that $L^{1},L^{2}$ agree on $I\times[0,T]$ if $L^{1}_{a,b}(x,y)=L^{2}_{a,b}(x,y)$ for all $0\leq a<b\leq T$ and all $a,b\in I.$ Just as we did in 3.2, we construct three copies $\mathcal{L}^{0},\mathcal{L}^{1},\mathcal{L}^{2}$ of the directed landscape, so that $\mathcal{L}^{1},\mathcal{L}^{2}$ are independent, and furthermore, with high probability, $\mathcal{L}^{0}$ agrees with $\mathcal{L}^{1}$ on $I_{r}\times[0,T]$ and $\mathcal{L}^{0}$ agrees with $\mathcal{L}^{2}$ on $J_{r}\times[0,T]$ for all $1\leq r\leq N$. We have the following construction based on the same argument as 3.2.
Consider fields of independent, identically distributed exponential random variables $\omega^{i}=\{\omega^{i}(t,x)\}_{t\geq 0,x\in\mathbb{Z}}$ for $i=1,2$, independent of one another. Define a third field of independent, identically distributed variables $\omega_{\varepsilon}^{0}=\{\omega_{\varepsilon}^{0}(t,x)\}_{t\geq 0,x\in\mathbb{Z}}$ which agrees with $\omega^{1}$ on $\mathbb{Z}_{\geq 0}\times\bigcup((\varepsilon^{-2/3}\bar{I}_{r})\cap\mathbb{Z})$ and agrees with $\omega^{2}$ on $\mathbb{Z}_{\geq 0}\times\bigcup((\varepsilon^{-2/3}\bar{J}_{r})\cap\mathbb{Z})$. Just as we did in 3.2, we take a joint limit point of the last passage times for all three fields (together with all of the geodesic paths associated to the finite collection of endpoints given by $b_{1},c_{1},...,b_{N},c_{N}$) as $\varepsilon\to 0$. We will obtain three copies of the directed landscape $\mathcal{L}^{0},\mathcal{L}^{1},\mathcal{L}^{2}$ coupled onto the same space (together with the respective geodesics).
Figure 1. A visual depiction of the “geodesic separation” argument in the proof of 3.4. We view the horizontal axis as a spatial axis, and likewise view the vertical axis as a temporal one. If all of the prelimiting geodesic paths from $b_{r}$ to $b_{r}$ and from $c_{r}$ to $c_{r}$ do not intersect the dotted lines or the endpoints of the intervals, then the non-crossing property of the geodesics in the prelimiting model implies that all last passage values between two space-time points inside the left shaded region will agree with those determined by $\omega^{1}$, while all last passage values between two space-time points inside the right shaded region will agree with those determined by $\omega^{2}$.
Consider the event $E$ in which all of the $4N$ geodesic paths from the points $b_{r}$ to $b_{r}$ and $c_{r}$ to $c_{r}$ in $\mathcal{L}^{0}$ do not cross any of the values $a_{r}$, and also do not cross the endpoints of the intervals $I_{r},J_{r}$ (which in particular implies that they do not intersect). By the same argument as in 3.2, the values $\mathcal{L}^{0}_{s,t}(x,y)$ will agree with either of $\mathcal{L}^{1}$ or $\mathcal{L}^{2}$ as long as $x,y$ lie strictly in between any given consecutive geodesics $\Pi_{r}$ and $\Pi_{r+1}$. See Fig. 1 for a geometric depiction of this event.
We apply a union bound over all $4N$ of the geodesics, together with an application of 3.1, to show that the event $E$ occurs with probability greater than $1-C(4N)e^{-dD^{3}}$, where $C,d$ do not depend on $N,D$ or the choice of intervals. Consequently, by repeating the exact argument as that in 3.3 (replacing the event $\{X>k\}$ appearing there with the event $E$ here) we obtain the required bound. ∎
Specializing 3.4 to the case $T=1$ and projecting the result to the two-time marginal $(s,t)=(0,1)$, we obtain the following corollary about strong mixing of the Airy sheet under diagonal shifts.
Corollary 3.5.
Consider the Airy sheet $\mathcal{S}=\mathcal{L}_{0,1}$. Fix $T>0$, and consider any compact disjoint intervals $I_{1},...,I_{N},J_{1},...,J_{N}\subset\mathbb{R}^{2}$. Define the $\sigma$-fields
$\mathcal{G}_{I}$ $:=\sigma(\{\mathcal{S}(x,y):(x,y)\in I_{r}\times I_{r}\text{ for some }r\leq N\}),$
$\mathcal{G}_{J}$ $:=\sigma(\{\mathcal{S}(x,y):(x,y)\in J_{r}\times J_{r}\text{ for some }r\leq N\}).$
Define
$D:=\inf\bigg{\{}|x-y|:x\in\bigcup_{1}^{N}I_{r},\;\;y\in\bigcup_{1}^{N}J_{r}\bigg{\}}.$
Then we have the bound
$|\operatorname{Cov}(F,G)|\leq CNe^{-dD^{3}}\mathbb{E}[F^{4}]^{1/4}\mathbb{E}[G^{4}]^{1/4}.$
Here $C,d>0$ are universal constants that are independent of $N,\{I_{r},J_{r}\}_{r=1}^{N},D$ and $F,G\in L^{4}(\Omega)$ that are respectively measurable with respect to $\mathcal{G}_{I}$ and $\mathcal{G}_{J}.$
4. Useful estimates on Airy sheets and argmaxes
We begin this section by establishing some useful notation.
Definition 4.1.
For $f,g\in C(\mathbb{R})$ and $L\in C(\mathbb{R}^{2})$ we define
$f\circ L\circ g:=\sup_{x,y\in\mathbb{R}}f(x)+L(x,y)+g(y).$
In this section, we prove a bound (4.9) on the covariance of $f\circ\mathcal{L}_{0,\epsilon}\circ g$ and $u\circ\mathcal{L}_{0,\epsilon}\circ v$ for deterministic functions $f,g,u,v$ all lying within a certain restricted domain. This estimate, and the additional estimates on the covariance of products of functions of this nature (4.10 and 4.11), are critical inputs to the proof of the estimate (2.5). Most of this section is devoted to setting up bounds which are used to prove 4.9, 4.10, and 4.11. The following lemma from [DSV22] is useful in our arguments.
Lemma 4.2 (Lemma 5.3, [DSV22]).
There exists a deterministic $c>0$ and a random $\mathfrak{C}>0$ so that
$|\mathcal{S}(x,y)+(x-y)^{2}|\leq\mathfrak{C}+c\log^{2/3}(2+|x|+|y|).$
Furthermore $\mathbb{E}[a^{\mathfrak{C}^{3/2}}]<\infty$ for some $a>1.$
We use this lemma to prove a helpful property of the point $(x,y)$ which maximizes $f\circ\mathcal{L}_{0,\epsilon}\circ g$.
Proposition 4.3.
Let $f,g:\mathbb{R}\to\mathbb{R}$ be deterministic continuous functions satisfying $-2x^{2}-B\leq f(x),g(x)\leq-\frac{1}{2}x^{2}+B$ for some $B>0.$ We define the random variable
$(X_{\epsilon},Y_{\epsilon}):=\mathrm{argmax}_{(x,y)\in\mathbb{R}^{2}}\;\;f(x)+\mathcal{L}_{0,\epsilon}(x,y)+g(y).$
We define $z_{\max}:=\mathrm{argmax}_{z\in\mathbb{R}}f(z)+g(z),$ and assume that this point is unique. Then for $M^{2}>1+100B$,
$\mathbb{P}(X_{\epsilon}^{2}+Y_{\epsilon}^{2}>M^{2})\leq Ce^{-d\epsilon^{-1/2}M^{3}}$
where $C,d>0$ are universal constants not depending on $f,g,M,\epsilon,B$.
Proof.
By the scale invariance of the directed landscape, we can replace $\mathcal{L}_{0,\epsilon}(x,y)$ by the process
$\epsilon^{1/3}\mathcal{L}_{0,1}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)=\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y).$
In the notation of 4.2 we have
$|\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)+\epsilon^{-1}(x-y)^{2}|\leq\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|x|+|y|)),$
where $c$ is deterministic and $\mathfrak{C}$ is a non-negative random variable satisfying $\mathbb{P}(\mathfrak{C}>u)\leq Ce^{-du^{3/2}}$ for $u\geq 0$. From now on, we use the notation $\mathcal{S}^{\epsilon}(x,y):=\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)$. In particular,
$\begin{split}(f+g)(z_{\max})-\mathfrak{C}\epsilon^{1/3}-c\epsilon^{1/3}\log^{2/3}(2+2\epsilon^{-2/3}|z_{\max}|)&\leq f(z_{\max})+\mathcal{S}^{\epsilon}(z_{\max},z_{\max})+g(z_{\max})\\ &\leq f(X_{\epsilon})+\mathcal{S}^{\epsilon}(X_{\epsilon},Y_{\epsilon})+g(Y_{\epsilon})\end{split}$ (4.1)
By the assumptions of the proposition, on the event $X_{\epsilon}^{2}+Y_{\epsilon}^{2}>M^{2}$ we have that
$f(X_{\epsilon})+g(Y_{\epsilon})\leq 2B-\frac{1}{2}(X_{\epsilon}^{2}+Y_{\epsilon}^{2})\leq 2B-\frac{1}{2}M^{2}.$
However, 4.2 also implies that
$\mathcal{S}^{\epsilon}(X_{\epsilon},Y_{\epsilon})\leq\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(\epsilon^{-2/3}(2+|X_{\epsilon}|+|Y_{\epsilon}|)).$
From the conditions imposed on $f,g$ by the statement of the proposition, it is clear that $z_{\max}\in[-2\sqrt{B},2\sqrt{B}]$. Combining all of this with (4.1) we obtain the following bound on $\mathfrak{C}$:
$2\mathfrak{C}\epsilon^{1/3}$ $\geq(f+g)(z_{\max})-2B+\frac{1}{2}M^{2}+c\epsilon^{1/3}\bigg{(}-\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))-\log^{2/3}(2+2\epsilon^{-2/3}|z_{\max}|)\bigg{)}$
$\geq-z_{\max}^{2}-4B+\frac{1}{2}M^{2}+c\epsilon^{1/3}\bigg{(}-\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))-\log^{2/3}(2+2\epsilon^{-2/3}|z_{\max}|)\bigg{)}$
$\geq-8B+\frac{1}{2}M^{2}+c\epsilon^{1/3}\bigg{(}-\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))-\log^{2/3}(2+2\epsilon^{-2/3}B)\bigg{)}.$
The deterministic constant $c$ may grow from line to line. Therefore, for sufficiently small $\epsilon>0$ (below some threshold that is deterministic and independent of $B\geq 2$), $c\epsilon^{1/3}\log^{2/3}(2+2\epsilon^{-2/3}B)\leq B$.
Under the additional assumption that $X_{\epsilon}^{2}+Y_{\epsilon}^{2}\leq(2M)^{2}$, we find that $c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))\leq\frac{1}{4}M^{2}$ for sufficiently small $\epsilon>0$ (below some threshold that is deterministic and independent of $M\geq 2$). Therefore, the last expression admits a lower bound by $-9B+\frac{1}{4}M^{2}.$ We have shown that
$\mathbb{P}(M^{2}\leq X_{\epsilon}^{2}+Y_{\epsilon}^{2}\leq(2M)^{2})\leq\mathbb{P}(2\mathfrak{C}\epsilon^{1/3}\geq-9B+\frac{1}{4}M^{2}),$
for $\epsilon\leq\epsilon_{0}$ with $\epsilon_{0}$ deterministic and independent of $B,M.$ Summing over dyadically sized concentric annuli, we obtain
$\mathbb{P}(X_{\epsilon}^{2}+Y_{\epsilon}^{2}>M^{2})$ $=\sum_{k=0}^{\infty}\mathbb{P}(4^{k}M^{2}<X_{\epsilon}^{2}+Y_{\epsilon}^{2}\leq 4^{k+1}M^{2})$
$\leq\sum_{k=0}^{\infty}\mathbb{P}(2\mathfrak{C}\epsilon^{1/3}\geq-9B+4^{k-1}M^{2})$
$\leq\sum_{k=0}^{\infty}Ce^{-d\epsilon^{-1/2}(4^{k-1}M^{2}-9B)^{3/2}}$
$\leq\sum_{k=0}^{\infty}Ce^{-d\epsilon^{-1/2}8^{k-2}(M^{2}-36B)^{3/2}}.$
In the last line we used the bound $9B\leq 9\cdot 4^{k}B$. We can obtain a factor of $8^{-k}$ in front of the exponential terms, through the bounds $e^{-u}\leq 2u^{-1}e^{-u/2}\leq 2u^{-1}e^{-u/8^{k}}$. This allows us to bound the infinite sum by a constant and obtain a bound only in terms of the first summand. In summary, the final line of the equation above can be bounded above by a quantity of the form
$C\epsilon^{1/2}(M^{2}-36B)^{-3/2}e^{-d\epsilon^{-1/2}(M^{2}-36B)^{3/2}},$
where the value of $d$ can be reduced if desired. If $M^{2}$ is larger than $1+100B$, then $(M^{2}-36B)^{-3/2}$ can be bounded above by $1$. Furthermore, $M^{2}>1+100B>72B$ implies that $M^{2}-36B>\frac{1}{2}M^{2},$ and therefore $e^{-d\epsilon^{-1/2}(M^{2}-36B)^{3/2}}\leq e^{-d\epsilon^{-1/2}M^{3}},$ where once again, we can make $c$ larger and $d$ smaller if desired. ∎
We use 4.3 to obtain the following bound on $|X_{\epsilon}-Y_{\epsilon}|$.
Proposition 4.4.
Let $f,g:\mathbb{R}\to\mathbb{R}$ be deterministic continuous functions satisfying $-2x^{2}-B\leq f(x),g(x)\leq-\frac{1}{2}x^{2}+B$ for some $B>0.$ Define the random variable
$(X_{\epsilon},Y_{\epsilon}):=\mathrm{argmax}_{(x,y)\in\mathbb{R}^{2}}\;\;f(x)+\mathcal{L}_{0,\epsilon}(x,y)+g(y).$
We define $z_{\max}:=\mathrm{argmax}_{z\in\mathbb{R}}f(z)+g(z),$ and assume that this point is unique. Then
$\mathbb{P}(|X_{\epsilon}-Y_{\epsilon}|>\epsilon^{\frac{1}{2}-\delta})\leq Ce^{-d\epsilon^{-1/2}(\epsilon^{-2\delta}-108B)^{3/2}}+Ce^{-d\log^{6/5}(1/\epsilon)},$
where $C,d>0$ are universal constants independent of $f,g,B,\epsilon\in(0,1].$
Intuitively, the bound $\epsilon^{\frac{1}{2}-\delta}$ should not be optimal, as one would expect based on the scaling properties of the landscape that the typical value of $|X_{\epsilon}-Y_{\epsilon}|$ should be of order $\epsilon^{2/3}$. Without additional regularity assumptions on $f$ or $g$, this is simply not true. However, with additional Hölder regularity assumptions we will later show that this is indeed true, see 4.6 below.
Proof.
As in previous arguments, we replace $\mathcal{L}_{0,\epsilon}(x,y)$ by $\epsilon^{1/3}\mathcal{L}_{0,1}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)=\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)$. From 4.2, we have the bound
$|\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)+\epsilon^{-1}(x-y)^{2}|\leq\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|x|+|y|)),$
where $c$ is deterministic and $\mathfrak{C}$ is a non-negative random variable satisfying $\mathbb{P}(\mathfrak{C}>u)\leq Ce^{-du^{3/2}}$ for $u\geq 0$. We estimate the probability of the event $K_{\epsilon}:=\{|X_{\epsilon}-Y_{\epsilon}|>\epsilon^{\frac{1}{2}-\delta}\}$. On one hand, on the event $K_{\epsilon}$, $\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}>\epsilon^{-2\delta}$. From the conditions on $f,g$ imposed by the statement of the proposition, it is also clear that $|z_{\max}|\leq\sqrt{2}B.$ Therefore, on $K_{\epsilon}$ we find that
$-\epsilon^{-2\delta}$ $+2B\geq-\epsilon^{-2\delta}+\sup_{x}f(x)+\sup_{y}g(y)$
$\geq f(X_{\epsilon})-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(Y_{\epsilon})$
$\geq\big{[}f(X_{\epsilon})+\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}X_{\epsilon},\epsilon^{-2/3}Y_{\epsilon})+g(Y_{\epsilon})\big{]}-\mathfrak{C}\epsilon^{1/3}-c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))$
$\geq\big{[}f(z_{\max})+\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}z_{\max},\epsilon^{-2/3}z_{\max})+g(z_{\max})\big{]}-\mathfrak{C}\epsilon^{1/3}-c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))$
$\geq f(z_{\max})+g(z_{\max})-\bigg{(}\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(2+2\epsilon^{-2/3}|z_{\max}|)\bigg{)}$
$\;\;\;\;\;-\mathfrak{C}\epsilon^{1/3}-c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))$
$\geq(-B-\frac{1}{2}z_{\max}^{2})+(-B-\frac{1}{2}z_{\max}^{2})-2\mathfrak{C}\epsilon^{\frac{1}{3}}$
$\;\;\;\;\;-c\epsilon^{1/3}\bigg{(}\log^{2/3}(2+2\epsilon^{2/3}|z_{\max}|)+\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))\bigg{)}$
$\geq-4B-2\mathfrak{C}\epsilon^{1/3}-c\epsilon^{1/3}\bigg{(}\log^{2/3}(2+2\epsilon^{-2/3}B)+\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))\bigg{)}.$
Since $c$ is deterministic, for sufficiently small deterministic $\epsilon>0$ (below some threshold independent of $B\geq 2$ say) we have an elementary bound $c\epsilon^{1/3}\log^{2/3}(2+2\epsilon^{-2/3}B)\leq 2B$. We define the event
$H_{\epsilon}:=\{|X_{\epsilon}|+|Y_{\epsilon}|\leq 1+\log^{2/5}(1/\epsilon)+100B\}.$
By 4.3 we find that $1-\mathbb{P}(H_{\epsilon})\leq Ce^{-d\log^{6/5}(1/\epsilon)}.$ On the event $H_{\epsilon}$ we likewise have that $c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))\leq 100B$, again assuming that $\epsilon$ is below some sufficiently small positive deterministic threshold. On $K_{\epsilon}\cap H_{\epsilon}$ we find that $2\mathfrak{C}\epsilon^{1/3}\geq\epsilon^{-2\delta}-108B.$ Therefore, we obtain
$\mathbb{P}(K_{\epsilon})$ $\leq\mathbb{P}(H_{\epsilon}\cap K_{\epsilon})+\mathbb{P}(H_{\epsilon}^{c})$
$\leq\mathbb{P}(2\mathfrak{C}\epsilon^{1/3}\geq\epsilon^{-2\delta}-108B)+Ce^{-d\log^{6/5}(1/\epsilon)}$
$\leq Ce^{-d\epsilon^{-1/2}(\epsilon^{-2\delta}-108B)^{3/2}}+Ce^{-d\log^{6/5}(1/\epsilon)}.$
This completes the proof. ∎
We introduce several quantities which will be important to our argument.
Definition 4.5.
Let $f,g:\mathbb{R}\to\mathbb{R}$ be continuous functions decaying parabolically at infinity, that is, $-Ax^{2}-B\leq f(x),g(x)\leq-Cx^{2}-D$ for some $A>C>0$ and $B,D\in\mathbb{R}.$ Assume that $f,g$ achieve their maximum values at unique input values $z_{\max}^{f},z_{\max}^{g}$. For $\epsilon,M>0$ and $\alpha\in(0,1)$, we define
$\mathscr{S}(f;\epsilon)$ :=$:=\{z\in\mathbb{R}:f(z)\geq f(z_{\max}^{f})-\epsilon\}$ (4.2)
$\Gamma(f,g;\epsilon)$ $:=\inf\{|z-z^{\prime}|:|f(z)-f(z_{\max}^{f})|<\epsilon,\;\;|g(z^{\prime})-g(z_{\max}^{g})|<\epsilon\}.$ (4.3)
$\mathfrak{Hol}(g,\alpha;M)$ $:=\sup_{\begin{subarray}{c}s,t\in[-M,M]\\ s\neq t\end{subarray}}\frac{|g(t)-g(s)|}{|t-s|^{\alpha}}.$
We define $N_{\epsilon}(f)$ to be the smallest number of intervals of length $\epsilon^{2}$ required to cover $\mathscr{S}(f;\epsilon).$
Proposition 4.6.
Let $f,g:\mathbb{R}\to\mathbb{R}$ be deterministic functions satisfying $-2x^{2}-B\leq f(x),g(x)\leq-\frac{1}{2}x^{2}+B$ for some $B\geq 1.$ Define the random variable
$(X_{\epsilon},Y_{\epsilon}):=\mathrm{argmax}_{(x,y)\in\mathbb{R}^{2}}\;\;f(x)+\mathcal{L}_{0,\epsilon}(x,y)+g(y).$
We also denote $z_{\max}:=\mathrm{argmax}_{z\in\mathbb{R}}f(z)+g(z),$ which we assume is unique. Fix $\delta\in(0,1/3]$. Then
[…]
where $C,d>0$ may depend on $\delta$ but are independent of $f,g,B$ and $\epsilon\in(0,1]$.
Here and in later propositions, we do not optimize the constants such as $3/10,2/5,6/5,100$. The current values suffice for arguments in Section 6.
Proof.
By scale invariance of the directed landscape we can replace $\mathcal{L}_{0,\epsilon}(x,y)$ by $\epsilon^{1/3}\mathcal{L}_{0,1}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)=\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)$, where $\mathcal{S}$ is the Airy sheet as usual. In the notation of 4.2 we have
$|\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)+\epsilon^{-1}(x-y)^{2}|\leq\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|x|+|y|)),$
where $c$ is deterministic and $\mathfrak{C}$ is a non-negative random variable satisfying $\mathbb{P}(\mathfrak{C}>u)\leq Ce^{-du^{3/2}}$ for $u\geq 0$. We choose $v:=\frac{3\delta}{10-6\delta}>\frac{3}{10}\delta$ and define an inductive sequence of exponents $\kappa_{n}$ for $n=0,1,2,3,...$ as follows:
$\kappa_{0}=\frac{1}{3}\text{ and }\kappa_{n+1}:=\frac{1-v}{2}+\frac{1}{2}\bigg{(}\frac{1}{2}-v\bigg{)}\kappa_{n}.$
As $n\to\infty$, $\kappa_{n}\uparrow\frac{2}{3}-\frac{1}{2}\delta$. This convergence necessarily occurs exponentially quickly by standard iteration theory for linear recurrences. For $n=0,1,2,3,...$ define the events
$K^{(n)}_{\epsilon}:=\{|X_{\epsilon}-Y_{\epsilon}|\leq\epsilon^{\kappa_{n}}\}.$
We will now prove that there exist universal constants $C,d>0$ so for all $n\geq 1$ and $\epsilon\in(0,1]$,
$\mathbb{P}(K^{(n-1)}_{\epsilon}\backslash K_{\epsilon}^{(n)})\leq 1_{\{B\geq d\epsilon^{-1}\}}+1_{\{\mathfrak{Hol}(g;\frac{1}{2}-\frac{v}{2};1+100B+\log^{2/5}(1/\epsilon))>\epsilon^{-\delta\kappa_{n-1}/2}\}}+Ce^{-d\log^{6/5}(1/\epsilon)}.$ (4.4)
To prove this, notice
$f(X_{\epsilon})$ $-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(Y_{\epsilon})$
$\geq f(X_{\epsilon})+\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}X_{\epsilon},\epsilon^{-2/3}Y_{\epsilon})+g(Y_{\epsilon})-\mathfrak{C}\epsilon^{1/3}-c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))$
$\geq f(z_{\max})+\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}z_{\max},\epsilon^{-2/3}z_{\max})+g(z_{\max})-\mathfrak{C}\epsilon^{1/3}-c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))$
$\geq f(z_{\max})+g(z_{\max})-\bigg{(}\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(2+2\epsilon^{-2/3}|z_{\max}|)\bigg{)}$
$\;\;\;\;\;\;-\mathfrak{C}\epsilon^{1/3}-c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))$
From the conditions on $f,g$ imposed by the statement of the proposition, it follows that $|z_{\max}|\leq\sqrt{2}B.$ Since $c$ is deterministic, for sufficiently small deterministic $\epsilon>0$, which does not depend on the choice of $B\geq 1$, we have an elementary bound $c\epsilon^{1/3}\log^{2/3}(2+2\epsilon^{-2/3}B)\leq C\epsilon^{\frac{1}{3}-\frac{2}{3}v}B^{v}.$ This follows from the fact that $\log^{2/3}(2+u)\leq Cu^{v}$ for all $u\geq 1$ where $C=C(v)$ does not depend on $u.$ We define $H_{\epsilon}$ to be the event
$H_{\epsilon}:=\{|X_{\epsilon}|+|Y_{\epsilon}|\leq 1+\log^{2/5}(1/\epsilon)+100B\}.$
By 4.3, we conclude $1-\mathbb{P}(H_{\epsilon})\leq Ce^{-d\log^{6/5}(1/\epsilon)}.$ Likewise, on the event $H_{\epsilon}$ we have $c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))\leq C\epsilon^{\frac{1}{3}-\frac{2}{3}v}B^{v}$. Here we have assumed again that $\epsilon$ is below some sufficiently small positive deterministic threshold. Consequently, on the event $H_{\epsilon}$, we obtain the following bound
$f(X_{\epsilon})-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(Y_{\epsilon})\geq f(z_{\max})+g(z_{\max})-2\mathfrak{C}\epsilon^{1/3}-C\epsilon^{\frac{1}{3}-\frac{2}{3}v}B^{v}.$ (4.5)
On the event $K^{(n-1)}_{\epsilon}\cap H_{\epsilon}$, $|g(X_{\epsilon})-g(Y_{\epsilon})|\leq C\epsilon^{-v\kappa_{n-1}/2}|X_{\epsilon}-Y_{\epsilon}|^{\frac{1}{2}-\frac{v}{2}}\leq\epsilon^{(\frac{1}{2}-v)\kappa_{n-1}}$ as long as we have $\mathfrak{Hol}(g;\frac{1}{2}-\frac{v}{2};1+100B+\log^{2/5}(1/\epsilon))\leq\epsilon^{-v\kappa_{n-1}/2}.$ When we combine this bound with (4.5), on the event $H_{\epsilon}\cap K^{(n-1)}_{\epsilon}$, we obtain the bound
$f(z_{\max})-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(z_{\max})$ $\geq f(X_{\epsilon})-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(X_{\epsilon})$
$\geq f(X_{\epsilon})-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(Y_{\epsilon})-|g(X_{\epsilon})-g(Y_{\epsilon})|$
$\geq f(z_{\max})+g(z_{\max})-2\mathfrak{C}\epsilon^{1/3}-C\epsilon^{\frac{1}{3}-\frac{2}{3}v}B^{v}-\epsilon^{(\frac{1}{2}-v)\kappa_{n-1}}$
$\geq f(z_{\max})+g(z_{\max})-2\mathfrak{C}\epsilon^{1/3}-C\epsilon^{(\frac{1}{2}-v)\kappa_{n-1}}B^{v}.$
In the final line, we use the fact that $B\geq 1$ and that $(\frac{1}{2}-v)\kappa_{n}$ is always less than $\frac{1}{3}-\frac{2}{3}v.$ As long as $\mathfrak{Hol}(g;\frac{1}{2}-\frac{v}{2};1+100B+\log^{2/5}(1/\epsilon))\leq\epsilon^{-v\kappa_{n-1}/2}$, we have by the recursive definition of the $\kappa_{n}$ that $1+(\frac{1}{2}-v)\kappa_{n-1}=2\kappa_{n}+v$ and therefore the above expression immediately gives
$(X_{\epsilon}-Y_{\epsilon})^{2}\leq 2\mathfrak{C}\epsilon^{4/3}+C\epsilon^{1+(\frac{1}{2}-v)\kappa_{n-1}}B^{v}=2\mathfrak{C}\epsilon^{4/3}+C\epsilon^{2\kappa_{n}+v}B^{v},$
for some large enough deterministic constant $C=C(v)$ that does not depend on $B,f,g,\epsilon.$ The right-hand side of the previous expression is less than $\epsilon^{2\kappa_{n}}$ as long as $\mathfrak{C}\leq\frac{1}{4}\epsilon^{-2v}$ (because $2\kappa_{n}\leq\frac{4}{3}-2v$ for all $n)$ and $B\leq\frac{1}{2}\epsilon^{-1}C^{-\frac{1}{v}},$ which proves (4.4) once we use the facts that $\mathbb{P}(H_{\epsilon}^{c})\leq Ce^{-d\log^{6/5}(1/\epsilon)}$ and $\mathbb{P}(\mathfrak{C}>\frac{1}{2}\epsilon^{-2v})\leq Ce^{-d\epsilon^{-3v}}\leq Ce^{-d\log^{6/5}(1/\epsilon)}$.
We use (4.4) to prove the bound in the proposition. By using 4.4 and the fact that $\kappa_{0}$ is chosen to be 1/3,
$\mathbb{P}(K^{(0)}_{\epsilon})\leq Ce^{-d\epsilon^{-1/2}(\epsilon^{-1/3}-108B)^{3/2}}\leq Ce^{-d\epsilon^{-5/6}}+1_{\{200B\geq\epsilon^{-1/3}\}}.$
Using this bound and then repeatedly summing over (4.4) up to a large enough value of $n$ such that $\kappa_{n}$ reaches the desired threshold will give the desired bound (this is a finite sum). This is because $v>\frac{3}{10}\delta$ for $\delta\in(0,1/3]$, as explained earlier. ∎
Proposition 4.7.
Let $f,g:\mathbb{R}\to\mathbb{R}$ be deterministic functions satisfying $-2x^{2}-B\leq f(x),g(x)\leq-\frac{1}{2}x^{2}+B$ for some $B\geq 1.$ Define the random variable
$(X_{\epsilon},Y_{\epsilon}):=\mathrm{argmax}_{(x,y)\in\mathbb{R}^{2}}\;\;f(x)+\mathcal{L}_{0,\epsilon}(x,y)+g(y).$
We also denote $z_{\max}:=\mathrm{argmax}_{z\in\mathbb{R}}f(z)+g(z)$, which we assume is unique. Fix $\delta\in(0,1/3]$. Then
[…]
where $C,d>0$ may depend on $\delta$ but are independent of $f,g,B$ and $\epsilon\in(0,1]$.
Proof.
The bound on $Y_{\epsilon}$ follows from the bound on $X_{\epsilon}$ by a symmetry argument. Therefore, we focus on the first bound. By the scale invariance of the directed landscape, we replace $\mathcal{L}_{0,\epsilon}(x,y)$ by $\epsilon^{1/3}\mathcal{L}_{0,1}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)=\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)$. In the notation of 4.2 we have
$|\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)+\epsilon^{-1}(x-y)^{2}|\leq\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|x|+|y|)),$
where $c$ is deterministic and $\mathfrak{C}$ is a non-negative random variable satisfying $\mathbb{P}(\mathfrak{C}>u)\leq Ce^{-du^{3/2}}$ for $u\geq 0$. We fix an arbitrary $\delta_{1}>0$ (the value of which will be specified at a later point). We define the event
$H_{\epsilon}:=\{|X_{\epsilon}|+|Y_{\epsilon}|\leq 1+\log^{2/5}(1/\epsilon)+100B\}.$
We recall on $H_{\epsilon}$, from (4.5) for $\epsilon$ below a sufficiently small positive deterministic threshold
$f(X_{\epsilon})-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(Y_{\epsilon})\geq f(z_{\max})+g(z_{\max})-2\mathfrak{C}\epsilon^{1/3}-C\epsilon^{\frac{1}{3}-\frac{2}{3}\delta_{1}}B^{\delta_{1}}.$
We also define the event
$F_{\epsilon}:=\{|X_{\epsilon}-Y_{\epsilon}|\leq\epsilon^{\frac{2}{3}-\delta_{1}}\}.$
By 4.6, $F_{\epsilon}$ has probability greater than $1-Ce^{-d\log^{6/5}(1/\epsilon)}$ as long as $B\leq d\epsilon^{-1/3}$ and
$\mathfrak{Hol}\big{(}g,\frac{1}{2}(1-\delta_{2}),1+100B+\log^{2/5}(1/\epsilon)\big{)}\leq\epsilon^{-\delta_{2}/3},$
where $\delta_{2}:=\frac{3}{10}\delta_{1}$. From the expression above, if $\mathfrak{Hol}\big{(}g,\frac{1}{2}(1-\delta_{2}),1+100B+\log^{2/5}(1/\epsilon)\big{)}\leq\epsilon^{-\delta_{2}/3}$, then on the event $F_{\epsilon}\cap H_{\epsilon}$ we have the bound
$f(X_{\epsilon})+g(X_{\epsilon})$ $>f(X_{\epsilon})-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(X_{\epsilon})$
$\geq f(X_{\epsilon})-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(Y_{\epsilon})-|g(Y_{\epsilon})-g(X_{\epsilon})|$
$\geq f(X_{\epsilon})-\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}+g(Y_{\epsilon})-\epsilon^{-\delta_{2}/3}|X_{\epsilon}-Y_{\epsilon}|^{\frac{1}{2}-\delta_{2}}$
$\geq f(z_{\max})+g(z_{\max})-2\mathfrak{C}\epsilon^{1/3}-C\epsilon^{\frac{1}{3}-\frac{2}{3}\delta_{1}}B^{\delta_{1}}-\epsilon^{\frac{1}{3}-\delta_{2}-\frac{1}{2}\delta_{1}+\delta_{1}\delta_{2}}.$
With $\delta$ as in the proposition statement, we make a specific choice of $\delta_{1}$ so that $\frac{2}{3}\delta_{1}<\frac{1}{2}\delta$ and $\delta_{2}+\frac{1}{2}\delta_{1}-\delta_{1}\delta_{2}<\frac{1}{2}\delta.$ For precision, take $\delta_{1}:=\frac{1}{2}\delta$. From the previous expression, it follows that in order to have $X_{\epsilon}\notin\mathscr{S}(f+g,\epsilon^{\frac{1}{3}-\delta})$, for $\epsilon<3^{-2/\delta}$ it must be true that $\mathfrak{C}>\frac{1}{6}\epsilon^{-\delta}$ or $B>\frac{1}{6}\epsilon^{-\frac{2}{3}}C^{-\frac{1}{\delta_{1}}}.$ Therefore,
$\mathbb{P}(X_{\epsilon}\notin\mathscr{S}(f+g,\epsilon^{\frac{1}{3}-\delta})\leq\mathbb{P}(\mathfrak{C}>\frac{1}{6}\epsilon^{-\delta})+1_{\{B>\frac{1}{6}\epsilon^{-\frac{2}{3}}C^{-\frac{1}{\delta_{1}}}\}}+\mathbb{P}(H_{\epsilon}^{c})+\mathbb{P}(F_{\epsilon}^{c}).$
To finish the proof, we apply the tail bounds on $\mathfrak{C}$ and the upper bounds on the latter two terms, all of which have already been discussed. ∎
Using the preceding bounds, we can prove the following estimate on the directed landscape.
Proposition 4.8.
Let $f,g:\mathbb{R}\to\mathbb{R}$ be deterministic functions satisfying $-2x^{2}-B\leq f(x),g(x)\leq-\frac{1}{2}x^{2}+B$ for some $B\geq 1.$ Then for any $\delta>0$ and $p>1$,
[…]
Here $C,d$ may depend on $p$ but are uniform over all $f,g,B,$ and $\epsilon\in(0,1]$.
Proof.
As before, by scale invariance of the directed landscape, we use the substitution, $\mathcal{L}_{0,\epsilon}(x,y)$ by $\epsilon^{1/3}\mathcal{L}_{0,1}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)=\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)$. In the notation of 4.2 we have
$|\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)+\epsilon^{-1}(x-y)^{2}|\leq\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|x|+|y|)),$ (4.6)
where $c$ is deterministic and $\mathfrak{C}$ is a non-negative random variable satisfying $\mathbb{P}(\mathfrak{C}>u)\leq Ce^{-du^{3/2}}$ for $u\geq 0$. Let $X_{\epsilon},Y_{\epsilon},z_{\max}$ be as in 4.6 and 4.7. We note that
$|f(X_{\epsilon})$ $+\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y)+g(Y_{\epsilon})-(f(z_{\max})+g(z_{\max}))|$
$\leq|f(X_{\epsilon})+g(X_{\epsilon})-(f(z_{\max})+g(z_{\max}))|+|\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}X_{\epsilon},\epsilon^{-2/3}Y_{\epsilon})+\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}|$
$\;\;\;\;\;+|g(Y_{\epsilon})-g(X_{\epsilon})|+\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}.$
We can now define the event
$J_{\epsilon}:=\{X_{\epsilon}\in\mathscr{S}(f+g,\epsilon^{\frac{1}{3}-\delta})\}\cap\{|X_{\epsilon}-Y_{\epsilon}|\leq\epsilon^{\frac{2}{3}-\delta}\}\cap\{|X_{\epsilon}|+|Y_{\epsilon}|\leq 1+100B+\log^{2/5}(1/\epsilon)\}.$
Furthermore, under the assumption that $\mathfrak{Hol}(g,\frac{1}{2}-\delta,1+100B+\log^{2/5}(1/\epsilon))<{\epsilon^{-\delta}},$ the last expression is bounded (term-by-term) by
$\epsilon^{\frac{1}{3}-\delta}+\big{(}\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|X_{\epsilon}|+|Y_{\epsilon}|))\big{)}+\epsilon^{-\delta/3}|X_{\epsilon}-Y_{\epsilon}|^{\frac{1}{2}-\delta}+\epsilon^{-1}(X_{\epsilon}-Y_{\epsilon})^{2}\\ \leq\epsilon^{\frac{1}{3}-\delta}+(\mathfrak{C}\epsilon^{1/3}+\epsilon^{\frac{1}{3}-\frac{2}{3}\delta}B^{\delta})+\epsilon^{(\frac{2}{3}-\delta)(\frac{1}{2}-\delta)}+\epsilon^{\frac{1}{3}-2\delta}.$
For the second term we used the following argument: since $c$ is deterministic, for sufficiently small deterministic $\epsilon>0$ which does not depend on the choice of $B\geq 1$, we have an elementary bound $c\epsilon^{1/3}\log^{2/3}(2+2\epsilon^{-2/3}B)\leq C\epsilon^{\frac{1}{3}-\frac{2}{3}\delta}B^{\delta}.$ This bound follows immediately from the fact that $\log^{2/3}(2+u)\leq Cu^{\delta}$ for all $u\geq 1$, where $C=C(\delta)$ does not depend on $u.$ Note that all powers of $\epsilon$ in the last expression are at worst $\frac{1}{3}-2\delta$. Since $\mathfrak{C}$ is the only random term in the last expression, these bounds imply that
$\mathbb{E}[|f\circ\mathcal{L}_{0,\epsilon}\circ g-\max(f+g)|^{p}\cdot 1_{J_{\epsilon}}]^{1/p}\leq CB^{\delta}\epsilon^{\frac{1}{3}-2\delta}.$
Replacing $\delta$ by $2\delta$ we obtain a bound of the same form as in the proposition statement. It remains to bound what occurs outside $J_{\epsilon}.$ Using 4.3 and 4.6 and 4.7, and the elementary inequality $(a+b)^{1/2}\leq a^{1/2}+b^{1/2}$, we obtain
[…]
By Cauchy-Schwarz and Minkowski, we furthermore have
$\mathbb{E}[|f\circ\mathcal{L}_{0,\epsilon}\circ g-\max(f+g)|^{p}\cdot 1_{J_{\epsilon}^{c}}]^{1/p}\leq\bigg{(}\mathbb{E}[|f\circ\mathcal{L}_{0,\epsilon}\circ g|^{2p}]^{1/2p}+\max(f+g)\bigg{)}\cdot\mathbb{P}(J_{\epsilon}^{c})^{1/2}.$
Clearly, $\max(f+g)\leq 2B.$ We will also show that $\mathbb{E}[|f\circ\mathcal{L}_{0,\epsilon}\circ g|^{2p}]^{1/2p}$ is bounded above by $B$ times some absolute constant. As before, we write $\mathcal{L}_{0,\epsilon}$ as $\mathcal{S}^{\epsilon}(x,y):=\epsilon^{1/3}\mathcal{S}(\epsilon^{-2/3}x,\epsilon^{-2/3}y).$ We apply (4.6) to obtain that
$f(x)+\mathcal{S}^{\epsilon}(x,y)+g(y)$ $\leq 2B-\frac{1}{2}(x^{2}+y^{2})+\mathfrak{C}\epsilon^{1/3}+c\epsilon^{1/3}\log^{2/3}(2+\epsilon^{-2/3}(|x|+|y|))$
$\leq 2B-\frac{1}{2}(x^{2}+y^{2})+\mathfrak{C}\epsilon^{1/3}+2+|x|+|y|,$
where the last bound is valid for sufficiently small $\epsilon>0$ below some deterministic threshold. Let $m$ denote the maximum value achieved by the function $(x,y)\mapsto 2+|x|+|y|-\frac{1}{2}(x^{2}+y^{2})$ over all of $\mathbb{R}^{2}$. Then in order for the last expression to exceed $2B+m+u$, it would be necessary for the constant $\mathfrak{C}$ to exceed $u$. Therefore,
$\mathbb{P}(f\circ\mathcal{L}_{0,\epsilon}\circ g>2B+m+u)\leq Ce^{-du^{3/2}},$ (4.8)
which implies that $\mathbb{E}[(f\circ\mathcal{L}_{0,\epsilon}\circ g)_{+}^{2p}]^{1/2p}$ (where $u_{+}:=\max\{u,0\}$) is bounded above by $B$ multiplied by an absolute constant. To bound the negative part, we note by similar arguments that the probability that $f\circ\mathcal{L}_{0,\epsilon}\circ g<-2B$ yields a large value of $\mathfrak{C}$. We can show this by considering what happens at $(z_{\max},z_{\max})$. This yields an analogous bound on $\mathbb{E}[(f\circ\mathcal{L}_{0,\epsilon}\circ g)_{-}^{2p}]^{1/2p}$ (where $u_{-}=\max\{0,-u\}$). ∎
We now come to the main result of this section, which we obtain by applying the mixing results of Section 3 with the estimates proved in this section. We begin by stating a covariance bound that will be crucial to understanding the mechanisms behind the black noise property.
Theorem 4.9.
Consider functions $f,g,u,v$ each lying in between $-B-2x^{2}$ and $B-\frac{1}{2}x^{2}$. Recall $\Gamma$ and $N_{\epsilon}$ as defined in (4.3). Then for any $\delta>0$,
$|\operatorname{Cov}(f\circ\mathcal{L}_{0,\epsilon}\circ g,u\circ\mathcal{L}_{0,\epsilon}\circ v)|$ $\leq CB^{2\delta}\min\big{\{}\epsilon^{\frac{2}{3}-\delta},\big{(}N_{\epsilon^{\frac{1}{3}-\delta}}(f+g)+N_{\epsilon^{\frac{1}{3}-\delta}}(u+v)\big{)}e^{-c\epsilon^{-2}\Gamma\big{(}f+g,u+v;\epsilon^{\frac{1}{3}-\delta}\big{)}^{3}}\big{\}}$
$\;\;\;\;\;+CB^{2}\cdot(\mathfrak{E}_{\epsilon}(f;B)+\mathfrak{E}_{\epsilon}(g;B)+\mathfrak{E}_{\epsilon}(u;B)+\mathfrak{E}_{\epsilon}(v;B)).$
Here $C,d$ are universal constants uniform over all $f,g,u,v$, $\epsilon,B$ and
[…]
Proof.
By Cauchy-Schwartz,
$|\operatorname{Cov}(f\circ\mathcal{L}_{0,\epsilon}\circ g,u\circ\mathcal{L}_{0,\epsilon}\circ v)|$ $=\big{|}\operatorname{Cov}\big{(}f\circ\mathcal{L}_{0,\epsilon}\circ g-\max(f+g),u\circ\mathcal{L}_{0,\epsilon}\circ v-\max(u+v)\big{)}\big{|}$
$\leq\mathbb{E}[(f\circ\mathcal{L}_{0,\epsilon}\circ g-\max(f+g))^{2}]^{1/2}\mathbb{E}[(u\circ\mathcal{L}_{0,\epsilon}\circ v-\max(u+v))^{2}]^{1/2}.$
[…]
We define the event $J_{\epsilon}^{f,g}:=\{X_{\epsilon}^{f,g},Y_{\epsilon}^{f,g}\in\mathscr{S}(f+g,\epsilon^{\frac{1}{3}-\delta})\}\cap\{|X_{\epsilon}^{f,g}-Y_{\epsilon}^{f,g}|\leq\epsilon^{\frac{2}{3}-\delta}\}\cap\{|X_{\epsilon}^{f,g}|+|Y_{\epsilon}^{f,g}|\leq 1+100B+\log^{2/5}(1/\epsilon)\}$, where $X_{\epsilon}^{f,g},Y_{\epsilon}^{f,g}$ are the argmaxes as defined in 4.6 and 4.7 respectively. We define the event $J^{u,v}_{\epsilon}$ in the same way, but with $f,g$ replaced by $u,v$. Finally, we define $J_{\epsilon}^{f,g,u,v}:=J^{f,g}_{\epsilon}\cap J^{u,v}_{\epsilon}$. We condition on the event $J_{\epsilon}^{f,g,u,v}$, and this allows us to use 3.4 to obtain a bound of the form $\big{(}N_{\epsilon^{\frac{1}{3}-\delta}}(f+g)+N_{\epsilon^{\frac{1}{3}-\delta}}(u+v)\big{)}e^{-c\epsilon^{-2}\Gamma\big{(}f+g,u+v;\epsilon^{\frac{1}{3}-\delta}\big{)}^{3}}.$ Indeed by setting $T:=\epsilon$ in 3.4, we see that the value of $D$ in that theorem corresponds to $\Gamma(f+g,u+v,\epsilon^{\frac{1}{3}-\delta}),$ and furthermore $N_{\epsilon^{\frac{1}{3}-\delta}}(f+g)+N_{\epsilon^{\frac{1}{3}-\delta}}(u+v)$ gives an upper bound for the value of $N$ which appears in that theorem. Consequently, we apply 3.4 to obtain the bound
$\bigg{|}\mathrm{Cov}\bigg{(}\big{(}f\circ\mathcal{L}_{0,\epsilon}\circ g\big{)}\cdot 1_{J_{\epsilon}^{f,g,u,v}},\big{(}u\circ\mathcal{L}_{0,\epsilon}\circ v\big{)}\cdot 1_{J_{\epsilon}^{f,g,u,v}}\bigg{)}\bigg{|}\leq\big{(}N_{\epsilon^{\frac{1}{3}-\delta}}(f+g)+N_{\epsilon^{\frac{1}{3}-\delta}}(u+v)\big{)}e^{-c\epsilon^{-2}\Gamma\big{(}f+g,u+v;\epsilon^{\frac{1}{3}-\delta}\big{)}^{3}},$
because on $J_{\epsilon}^{f,g,u,v}$ the maximum on all of $\mathbb{R}^{2}$ in the definitions of $f\circ\mathcal{L}_{0,\epsilon}\circ g,u\circ\mathcal{L}_{0,\epsilon}\circ v$ can be replaced by a maximum on some set of the form which appears in 3.4. It remains to control what happens on the complement of $J_{\epsilon}^{f,g,u,v}.$ By applying Cauchy-Schwarz twice we have
$\bigg{|}\mathrm{Cov}$ $\big{(}f\circ\mathcal{L}_{0,\epsilon}\circ g,u\circ\mathcal{L}_{0,\epsilon}\circ v\big{)}\bigg{|}\leq\bigg{|}\mathrm{Cov}\bigg{(}\big{(}f\circ\mathcal{L}_{0,\epsilon}\circ g\big{)}\cdot 1_{J_{\epsilon}^{f,g,u,v}},\big{(}u\circ\mathcal{L}_{0,\epsilon}\circ v\big{)}\cdot 1_{J_{\epsilon}^{f,g,u,v}}\bigg{)}\bigg{|}$
$\;\;\;\;+2\mathbb{E}[(f\circ\mathcal{L}_{0,\epsilon}\circ g-\max(f+g))^{4}]^{1/4}\mathbb{E}[(u\circ\mathcal{L}_{0,\epsilon}\circ v-\max(u+v))^{4}]^{1/4}\cdot(1-\mathbb{P}(J_{\epsilon}^{f,g,u,v}))^{\frac{1}{2}}.$
Using arguments similar to (4.7), it follows that $(1-\mathbb{P}(J_{\epsilon}^{f,g,u,v}))^{\frac{1}{2}}$ yields an upper bound consisting of terms of the form appearing in $\mathfrak{E}_{\epsilon}$. The upper bounds for all other terms have already been discussed. ∎
Theorem 4.10.
Consider functions $f,g,u,v,\bar{u},\bar{v}$ each lying in between $-B-2x^{2}$ and $B-\frac{1}{2}x^{2}$. Define $\Gamma(\epsilon):=\min\{\Gamma\big{(}f+g,u+v;\epsilon^{\frac{1}{3}-\delta_{1}}\big{)},\Gamma\big{(}f+g,\bar{u}+\bar{v};\epsilon^{\frac{1}{3}-\delta_{1}}\big{)}\}.$ Then for any $\delta>0$
$\bigg{|}\mathrm{Cov}\bigg{(}f\circ$ $\mathcal{L}_{0,\epsilon}\circ g\;\;,\;\;(u\circ\mathcal{L}_{0,\epsilon}\circ v-\max(u+v))((\bar{u}\circ\mathcal{L}_{0,\epsilon}\circ\bar{v}-\max(\bar{u}+\bar{v}))\bigg{)}\bigg{|}$
$\leq CB^{2\delta}\min\big{\{}\epsilon^{1-\delta},\big{(}N_{\epsilon^{\frac{1}{3}-\delta}}(f+g)+N_{\epsilon^{\frac{1}{3}-\delta}}(u+v)+N_{\epsilon^{\frac{1}{3}-\delta}}(\bar{u}+\bar{v})\big{)}e^{-c\epsilon^{-2}\Gamma(\epsilon)^{3}}\big{\}}$
$\;\;\;\;\;\;\;\;\;\;+CB^{3}\cdot(\mathfrak{E}_{\epsilon}(f;B)+\mathfrak{E}_{\epsilon}(g;B)+\mathfrak{E}_{\epsilon}(u;B)+\mathfrak{E}_{\epsilon}(v;B)+\mathfrak{E}_{\epsilon}(\bar{u};B)+\mathfrak{E}_{\epsilon}(\bar{v};B)).$
Here $C,d$ are universal constants uniform over all $f,g,u,v$, $\epsilon,B$, and $\mathfrak{E}_{\epsilon}$ was defined in 4.9.
Proof.
The proof is very similar to that of the previous theorem. Note, by applying Cauchy-Schwartz twice, that
$\bigg{|}\mathrm{Cov}\bigg{(}f\circ\mathcal{L}_{0,\epsilon}\circ g,(u\circ\mathcal{L}_{0,\epsilon}\circ v-\max(u+v))((\bar{u}\circ\mathcal{L}_{0,\epsilon}\circ\bar{v}-\max(\bar{u}+\bar{v}))\bigg{)}\bigg{|}\\ \leq\mathbb{E}[(f\circ\mathcal{L}_{0,\epsilon}\circ g-\max(f+g))^{2}]^{1/2}\mathbb{E}[(u\circ\mathcal{L}_{0,\epsilon}\circ v-\max(u+v))^{4}]^{1/4}\mathbb{E}[(\bar{u}\circ\mathcal{L}_{0,\epsilon}\circ\bar{v}-\max(\bar{u}+\bar{v}))^{4}]^{1/4}.$
[…]
Next, we proceed exactly as in the previous theorem. We condition on the event $J_{\epsilon}^{f,g,u,v,\bar{u},\bar{v}}$, which allows us to use 3.4 to obtain a bound of the form
$\big{(}N_{\epsilon^{\frac{1}{3}-\delta_{1}}}(f+g)+N_{\epsilon^{\frac{1}{3}-\delta_{1}}}(u+v)+N_{\epsilon^{\frac{1}{3}-\delta_{1}}}(\bar{u}+\bar{v})\big{)}e^{-c\epsilon^{-2}\Gamma(\epsilon)^{3}}.$
Indeed by setting $T:=\epsilon$ in that theorem, the value of $D$ corresponds to $\Gamma(\epsilon),$ and moreover $N_{\epsilon^{\frac{1}{3}-\delta_{1}}}(f+g)+N_{\epsilon^{\frac{1}{3}-\delta_{1}}}(u+v)+N_{\epsilon^{\frac{1}{3}-\delta_{1}}}(\bar{u}+\bar{v})$ gives an upper bound for the value of $N$ appearing in that theorem. It remains to control what occurs on the complement of $J_{\epsilon}^{f,g,u,v,\bar{u},\bar{v}}$. This can be done completely analogously to the proof of Theorem 4.9. As we observed there, this will only yield error terms of the form $\mathfrak{E}_{\epsilon}$. ∎
Proposition 4.11.
Consider functions $f_{i},g_{i},u_{j},v_{j}$ ($1\leq i\leq k,1\leq j\leq\ell)$ each lying in between $-B-2x^{2}$ and $B-\frac{1}{2}x^{2}$. Then for any $\delta_{1}>0$ there exists $\delta_{2}>0$ so that
$\bigg{|}\mathrm{Cov}\bigg{(}\prod_{i=1}^{k}(f_{i}\circ\mathcal{L}_{0,\epsilon}\circ g_{i}-\max(f_{i}+g_{i})),\prod_{j=1}^{\ell}(u_{j}\circ\mathcal{L}_{0,\epsilon}\circ v_{j}-\max(u_{j}+v_{j}))\bigg{)}\bigg{|}\\ \leq CB^{\delta_{2}}\epsilon^{\frac{1}{3}(k+\ell)-\delta_{1}}+CB^{k+\ell}\cdot(\sum_{i}\mathfrak{E}_{\epsilon}(f_{i};B)+\mathfrak{E}_{\epsilon}(g_{i};B)+\sum_{j}\mathfrak{E}_{\epsilon}(u_{j};B)+\mathfrak{E}_{\epsilon}(v_{j};B)).$
Here $C,d$ are universal constants uniform over all $f_{i},g_{i},u_{i},v_{i}$, $\epsilon,B$, and $\mathfrak{E}_{\epsilon}$ is as defined in 4.9.
Proof.
We apply Cauchy-Schwartz to the covariance, and then apply Hölder’s inequality to obtain
$\bigg{|}\mathrm{Cov}$ $\bigg{(}\prod_{i=1}^{k}(f_{i}\circ\mathcal{L}_{0,\epsilon}\circ g_{i}-\max(f_{i}+g_{i})),\prod_{j=1}^{\ell}(u_{j}\circ\mathcal{L}_{0,\epsilon}\circ v_{j}-\max(u_{j}+v_{j}))\bigg{)}\bigg{|}$
$\leq\mathbb{E}\bigg{[}\prod_{i=1}^{k}(f_{i}\circ\mathcal{L}_{0,\epsilon}\circ g_{i}-\max(f_{i}+g_{i}))^{2}\bigg{]}^{1/2}\mathbb{E}\bigg{[}\prod_{j=1}^{\ell}(u_{j}\circ\mathcal{L}_{0,\epsilon}\circ v_{j}-\max(u_{j}+v_{j}))^{2}\bigg{]}^{1/2}$
$\leq\prod_{i=1}^{k}\mathbb{E}\bigg{[}(f_{i}\circ\mathcal{L}_{0,\epsilon}\circ g_{i}-\max(f_{i}+g_{i}))^{2k}\bigg{]}^{1/(2k)}\prod_{j=1}^{\ell}\mathbb{E}\bigg{[}(u_{j}\circ\mathcal{L}_{0,\epsilon}\circ v_{j}-\max(u_{j}+v_{j}))^{2k}\bigg{]}^{1/(2k)}.$
We use 4.8 to obtain the claim (note here that unlike the previous two theorems, we do not need 3.4 as we are simply using the unrefined bound without leveraging the mixing property, as this will be sufficient later). ∎
Remark 4.12.
In all propositions above, the constants $1/2$ and $2$ in the parabolic decay rates of $f,g$ could have been replaced by $A^{-1}$ and $A$ for arbitrary $A>1$. This would affect the constants such as $100$ and $108$ as well as the constants $C,d$. While we cannot control the behavior of these constants as $A\uparrow\infty$, there is never an issue for a fixed finite $A$, no matter how large. This will be important later due to the nature of 2.6, because we will want $f,g$ to be “typical” sample paths of parabolic Airy2 processes of scales within the interval $[\eta,1]$ for some fixed but very small $\eta>0$. In this case $A$ can be as large as $C\eta^{-1}$ for some universal $C>0.$
[…]
5. Estimates on 3-dimensional Bessel processes
Lemma 5.1.
Let $R$ be a three-dimensional Bessel process, and let $p\in\mathbb{N}$. Then there exists $C_{p}>0$ such that we have a uniform bound over all $s_{1}\leq...\leq s_{p}$ and $\epsilon\in(0,1]$
$\mathbb{P}\bigg{(}\inf_{t\in[s_{i},s_{i}+1]}R(t)\leq\epsilon,\;\forall 1\leq i\leq p\bigg{)}ds_{1}\cdots ds_{p}\leq C_{p}\epsilon^{p}\prod_{i=0}^{p-1}\min\big{\{}1,(s_{i+1}-s_{i})^{-3/2}(1+\log_{+}^{3/2}(s_{i+1}-s_{i}))\big{\}},$
where $s_{0}:=0$.
Proof.
We will prove the bound first in the case that $\epsilon=1$. We interpret the Bessel process as the Euclidean norm of a three-dimensional standard Brownian motion, that is, $R(t)=|\vec{B}(t)|.$ It is a helpful fact that if $\vec{B}=(B_{1},B_{2},B_{3})$, then through a union bound and well-known formulas for the distribution of the supremum of a Brownian motion on $[0,1]$,
$\mathbb{P}(\sup_{t\in[0,1]}|\vec{B}(t)|>a)\leq 3\mathbb{P}(\sup_{t\in[0,1]}B_{1}(t)>a/3)\leq 6e^{-a^{2}/18}.$ (5.1)
It is also helpful to note the following fact
$\mathbb{P}\bigg{(}|\vec{B}(1)|>b,\;\inf_{t\in[0,1]}|\vec{B}(t)|\leq 1\bigg{|}|\vec{B}(0)|=a\bigg{)}\leq Ce^{-\frac{1}{36}(a^{2}+b^{2})},$ (5.2)
where $C>0$ does not depend on $a,b>0.$ This follows from the strong Markov property. Indeed, define the stopping time $\tau$ as the first time $t$ that $|\vec{B}(t)|=1$ and the Brownian motion $\vec{W}(t):=\vec{B}(t+\tau)-\vec{B}(\tau)$. The event $\{|\vec{B}(1)|>b,\;\inf_{t\in[0,1]}|\vec{B}(t)|\leq 1\}$ is contained in the event $\{\sup_{t\in[0,1]}|\vec{W}(t)|>b-1\},$ and the second event is moreover independent of the $\mathcal{F}^{B}_{\tau}$-measurable event $\{\inf_{t\in[0,1]}|\vec{B}(t)|<1\}=\{\tau\leq 1\}.$ Therefore, by conditioning on the latter event, we see that the left side of (5.2) is bounded above by
$\mathbb{P}(\sup_{t\in[0,1]}|\vec{W}(t)|>b-1)\cdot\mathbb{P}(\inf_{t\in[0,1]}|\vec{B}(t)|<1|\vec{B}(0)=a).$
By (5.1) this is bounded above by $36e^{-\frac{1}{18}((a-1)^{2}+(b-1)^{2})}.$ Now use $(b-1)^{2}\geq\frac{1}{2}b^{2}-1^{2}$ and $(a-1)^{2}\geq\frac{1}{2}a^{2}-1^{2}$ and we finally arrive at (5.2) with $C=36e^{\frac{1}{9}}$.
We return to proving the lemma in the case $\epsilon=1$. Without loss of generality, we will assume that $s_{i+1}-s_{i}>3$. In order for $\inf_{t\in[s_{i},s_{i}+1]}R(t)\leq 1$ for all $11\leq i\leq p+$ 1, at least one of the following three things must happen for each $i=1,...,p$:
  1. (1)
    $\sup_{t\in[s_{i},s_{i}+1]}|B(t)-B(s_{i}+1)|>9\sqrt{\log(s_{i}-s_{i-1})}$.
  2. (2)
    $|B(s_{i+1})-B(s_{i}+1)|\leq 18\sqrt{\log(s_{i}-s_{i-1})}.$
  3. (3)
    $\sup_{t\in[s_{i+1},s_{i+1}+1]}|B(t)-B(s_{i+1})|>9\sqrt{\log(s_{i}-s_{i-1})}$.
We partition $[s_{1},s_{p})$ into $2p$ disjoint intervals of the form $[0,s_{1}),[s_{1},s_{1}+1),[s_{1}+1,s_{2}),[s_{2},s_{2}+1),[s_{2}+1,s_{3}),...,[s_{p},s_{p}+1),$ which we denote by $I_{1},J_{1},I_{2},J_{2},I_{3},...,J_{p}$. For a given realization of the Brownian motion $\vec{B}$, we say that an interval $I_{i}$ or $J_{i}$ is active if one of the conditions above holds on that interval. We have shown that among any three consecutive sub-intervals in the partition $I_{1},J_{1},I_{2},J_{2},I_{3},...,J_{p}$, at least one must be active. Some intervals may furthermore be doubly active in the sense that the conditions within the above conditions may be satisfied for two distinct values of $i$. Furthermore, at least one of the first two sub-intervals $I_{1},J_{1}$ must be active. Consequently, by counting the doubly active intervals twice, there must be at least $p$ sub-intervals from $I_{1},J_{1},I_{2},J_{2},I_{3},...,J_{p}$ that are active, each corresponding to a unique value of $i$ in the list above. Notice that $e^{-(9\sqrt{\log(s_{i+1}-s_{i})})^{2}/18}\leq 1\wedge(s_{i+1}-s_{i})^{-2}\leq 1\wedge(s_{i+1}-s_{i})^{-3/2}$, and the standard $3$-dimensional Gaussian tail bound tells us that
$\mathbb{P}\big{(}|B(s_{i+1})-B(s_{i}+1)|\leq 18\sqrt{\log(s_{i+1}-s_{i})}\big{)}\leq\max\{1,18^{3}\cdot(s_{i+1}-s_{i})^{-3/2}(1+\log^{3/2}_{+}(s_{i+1}-s_{i}))\}.$ (5.3)
By applying a union bound over all possible choices of $p$-tuples of active and doubly active sub-intervals from $I_{1},J_{1},I_{2},J_{2},I_{3},...,J_{p}$, and applying the independence of increments of the $3$-dimensional Brownian motion, we obtain the bound in the lemma. This is a consequence of (5.2) for doubly active intervals and (5.1) or (5.3) for active ones. Each of the bounds for the increments $[s_{i+1},s_{i}]$ will be used exactly once. This proves the claim when $\epsilon=1$.
For arbitrary $\epsilon>0$, we need to explain why we obtain a factor $\epsilon^{p}$. We begin by explaining this factor in the case that $p=1$. If we condition on reaching the value $1$ inside the interval $[s_{i},s_{i}+1]$, the probability of reaching $\epsilon$ inside that interval is bounded by the probability of reaching $\epsilon$ ever again after time $s_{i}$. Using the Markov property, and the optional stopping theorem for the martingale $R^{-1}$ (stopping when this martingale hits $\epsilon^{-1})$, we show that the latter probability is exactly $\epsilon.$ Therefore, for the $p=1$ case we have the bound
$\mathbb{P}\big{(}\inf_{t\in[s_{i},s_{i}+1]}R(t)<\epsilon\big{)}$ $=\mathbb{P}\bigg{(}\inf_{t\in[s_{i},s_{i}+1]}R(t)<\epsilon\bigg{|}\inf_{t\in[s_{i},s_{i}+1]}R(t)<1\bigg{)}\mathbb{P}\big{(}\inf_{t\in[s_{i},s_{i}+1]}R(t)<1\big{)}$
$\leq\epsilon\;\;\cdot\;\;s_{1}^{-3/2}(1+\log_{+}^{3/2}s_{1}).$
While this illustrates the bound in the case $p=1$, the same argument extends in a straightforward way to general values of $p$. ∎
Proposition 5.2.
Let $R$ be a standard two-sided $3$-dimensional Bessel process. Let $N_{\epsilon}(R)$ denote the smallest number of intervals of length $\epsilon^{2}$ needed to cover the random set $\{t\in\mathbb{R}:R(t)\leq\epsilon\}.$ Then $\sup_{\epsilon\in(0,1]}\mathbb{E}[N_{\epsilon}(R)^{p}]<\infty$ for all $p\geq 1.$
Proof.
This random variable is bounded above by
[…]
By scale invariance of the Bessel process, we know that the above random variable has the same law as
[…]
Let $p$ be an even integer, and take the $p^{th}$ moment, then apply Fubini to interchange the integral and expectation. We obtain that
[…]
By 5.1 the integrand is bounded above by
$C_{p}\prod_{i=0}^{p-1}\min\big{\{}1,(s_{i+1}-s_{i})^{-3/2}(1+\log_{+}^{3/2}(s_{i+1}-s_{i}))\big{\}},$
where $s_{0}:=0$. The fact that previous integral is finite follow from this bound. ∎
Definition 5.3.
We define a Bessel process pinned at $t\in\mathbb{R}$ to be the process $R(\bullet-t)$ where $R$ is a two-sided $3$-dimensional Bessel process, centered at zero. We will denote by $\mathbb{P}_{\mathrm{Bes3}}^{t_{1},t_{2}}$ to be the law on the canonical space $C(\mathbb{R})^{2},$ of two independent Bessel processes $R_{1},R_{2}$ pinned at $t_{1},t_{2}$ respectively.
Proposition 5.4 (The key estimate).
Fix $\delta\in(0,1/2)$ and $10<\bar{\gamma}<\gamma<$ 1. Then there exists $C>0$ so that for all $\epsilon\in(0,1]$ we have the estimate
$\sup_{|t_{1}-t_{2}|\geq\epsilon^{-\gamma}}$ $\mathbb{P}_{\mathrm{Bes3}}^{t_{1},t_{2}}\bigg{(}\mathrm{there\;\;exist\;\;}s_{1},s_{2}\in\mathbb{R}\mathrm{\;\;such\;\;that\;\;}R_{1}(s_{1})+R_{2}(s_{2})\leq\epsilon^{-\delta}\mathrm{\;\;and\;\;}|s_{1}-s_{2}|\leq 2\epsilon^{-\bar{\gamma}}\bigg{)}$
$\leq C\epsilon^{\frac{3}{2}\gamma-\delta(2+\gamma-\bar{\gamma})-\frac{1}{2}\bar{\gamma}}.$
For the black noise result, we will see that this bound is most useful when $\gamma$ is slightly smaller than $1/3$ and $\delta,\bar{\gamma}$ are very close to 0. In this case, note that the right side can be made arbitrarily close to $\epsilon^{1/2}$. In practice, we will only use this estimate with $\gamma=4/15$ and $\bar{\gamma}=\delta$ close to 0.
Proof.
The proof will proceed using a simple union bound. If such values of $s_{1},s_{2}\in\mathbb{R}$ exist as specified above, then there exists some $n\in\mathbb{Z}$ such that all of the following simultaneously occur: $|s_{1}-n|\leq\epsilon^{-\bar{\gamma}}$, $|s_{s}-n|\leq\epsilon^{-\bar{\gamma}}$, $R_{1}(s_{1})\leq\epsilon^{-\delta},R_{2}(s_{2})\leq\epsilon^{-\delta}.$ Applying this union bound over all $n\in\mathbb{Z}$, we find that the desired probability is upper bounded by
$\sum_{n\in\mathbb{Z}}\mathbb{P}_{\mathrm{Bes3}}^{t_{1},t_{2}}\bigg{(}\text{ there exist }s_{1},s_{2}\in\mathbb{R}\text{ such that }|s_{1}-\epsilon^{-\bar{\gamma}}n|\leq\epsilon^{-\bar{\gamma}},|s_{2}-\epsilon^{-\bar{\gamma}}n|\leq\epsilon^{-\bar{\gamma}},R_{1}(s_{1})\leq\epsilon^{-\delta},R_{2}(s_{2})\leq\epsilon^{-\delta}\bigg{)}.$
Without any loss of generality we can replace $(t_{1},t_{2})$ by $(0,t_{2}-t_{1})$ and assume that $T:=t_{2}-t_{1}>0,$ thus by assumption $T>\epsilon^{-\gamma}$. We can estimate the probability in the summand, using the scaling invariance of the Bessel process to rewrite
$\mathbb{P}_{\mathrm{Bes3}}^{0,T}\bigg{(}\text{ there exist }s_{1},s_{2}\in\mathbb{R}\text{ such that }|s_{1}-\epsilon^{-\bar{\gamma}}n|\leq\epsilon^{-\bar{\gamma}},|s_{2}-\epsilon^{-\bar{\gamma}}n|\leq\epsilon^{-\bar{\gamma}},R_{1}(s_{1})\leq\epsilon^{-\delta},R_{2}(s_{2})\leq\epsilon^{-\delta}\bigg{)}$
$=\mathbb{P}_{\mathrm{Bes3}}^{0,\epsilon^{\bar{\gamma}}T}\bigg{(}\text{ there exist }s_{1},s_{2}\in\mathbb{R}\text{ such that }|s_{1}-n|\leq 1,|s_{2}-n|\leq 1,R_{1}(s_{1})\leq\epsilon^{\frac{1}{2}\bar{\gamma}-\delta},R_{2}(s_{2})\leq\epsilon^{\frac{1}{2}\bar{\gamma}-\delta}\bigg{)}$
$=\prod_{i=1}^{2}\mathbb{P}_{\mathrm{Bes3}}^{0,\epsilon^{\bar{\gamma}}T}\bigg{(}\text{ there exists }s_{i}\in[n-1,n+1]\text{ such that }R_{i}(s_{i})\leq\epsilon^{\frac{1}{2}\bar{\gamma}-\delta}\bigg{)}$
$\leq C\bigg{[}\epsilon^{\frac{1}{2}\bar{\gamma}-\delta}|n|^{-\frac{3}{2}+\delta}\bigg{]}\bigg{[}\epsilon^{\frac{1}{2}\bar{\gamma}-\delta}|\epsilon^{\bar{\gamma}}T-n|^{-\frac{3}{2}+\delta}\bigg{]}$
$=C\epsilon^{\bar{\gamma}-2\delta}|n|^{-\frac{3}{2}+\delta}|\epsilon^{\bar{\gamma}}T-n|^{-\frac{3}{2}+\delta}.$
In the third line, we used the independence of $R_{1}$ and $R_{2}$, and 5.1 with $p=1$ in the fourth line (we replaced the logarithmic correction by a weaker bound of size $|n|^{\delta}$ for simplicity). Now notice that uniformly over $N\in\mathbb{N}$ we have by symmetry
$\sum_{n=1}^{N}\frac{1}{n^{\frac{3}{2}-\delta}(N-n)^{\frac{3}{2}-\delta}}\leq 2\sum_{1\leq n\leq N/2}\frac{1}{n^{\frac{3}{2}-\delta}(N-n)^{\frac{3}{2}-\delta}}\leq 2\left(\frac{2}{N}\right)^{\frac{3}{2}-\delta}\sum_{1\leq n\leq N/2}n^{-\frac{3}{2}+\delta}\leq\frac{2^{5/2}\zeta(\tfrac{3}{2}-\delta)}{N^{\frac{3}{2}-\delta}}.$
Using this bound with $N=\epsilon^{\bar{\gamma}}T\geq\epsilon^{\bar{\gamma}-\gamma},$ we find that the desired probability is upper bounded by $C\epsilon^{\bar{\gamma}-2\delta}(\epsilon^{\bar{\gamma}}T)^{-\frac{3}{2}+\delta}\leq C\epsilon^{\bar{\gamma}-2\delta}(\epsilon^{\bar{\gamma}-\gamma})^{-\frac{3}{2}+\delta}=C\epsilon^{\frac{1}{2}(3\gamma-\bar{\gamma})-\delta(2+\gamma-\bar{\gamma})}.$
6. Proof of the main theorem
Proposition 6.1.
Consider a standard Brownian motion $B$ on $[0,1]$, and let $T$ denote its unique argmax. Fix $\alpha\in(0,1)$. Define the process $U:=(B(T)-B(T+t))_{t\in[-\alpha T,\alpha(1-T)]}$. Consider a three-dimensional Bessel process $R$ independent of $T$ and $B$, and define the process $V:=(R(t))_{t\in[-\alpha T,\alpha(1-T)]}$. Then $U$ is absolutely continuous with respect to $V$, and the Radon-Nikodym derivative is deterministically bounded by $C(1-\alpha)^{-1/2}$ for some universal constant $C>0$.
In the statement of 6.1, we can think of a continuous function on a random closed interval $I\subset[-1,1]$ as an element of $C[-1,1]$ that is constant away from $I$. The last statement is sharp: if $\alpha=1$ then the Radon-Nikodym is only in $L^{3-\delta}$ as opposed to $L^{\infty}$, as discussed in [DSV22, Theorem 2.15].
Proof.
Conditioned on $T$, the process $(B(T)-B(T+t))_{t\in[-T,1-T]}$ is a two-sided Brownian meander, as discussed in [DSV22, Theorem 2.15]. The left side is a meander of length $T$ and the right side is a meander of length $1-T$. When we condition on $T$, these two meanders are independent of one another. Therefore, it suffices to show that for a Brownian meander $X$ of length $1$, the Radon-Nikodym derivative $(X(t))_{t\in[0,\alpha]}$ is absolutely continuous with respect to a Bessel process on the same interval $(R(t))_{t\in[0,\alpha]}.$
In the proof of [DSV22, Theorem 2.15], the authors observe that if $\alpha=1$, then the Radon-Nikodym derivative is simply given by $\sqrt{\pi/2}\cdot R(1)^{-1}.$ Let $\mathcal{F}^{R}_{t}$ denote the filtration of the Bessel process. Then for general $\alpha\in[0,1]$ the Radon-Nikodym derivative of $X$ with respect to $R$ on the interval $[0,\alpha]$ is simply given by
$\sqrt{\pi/2}\cdot\mathbb{E}[R(1)^{-1}|\mathcal{F}^{R}_{\alpha}]=\sqrt{\pi/2}\cdot f_{\alpha}(R(\alpha)),$
where the expectation is taken with respect to a $3$-dimensional Bessel Process and $f_{\alpha}:[0,\infty)\to(0,\infty)$ is the unique function satisfying
[…]
In other words, $\vec{x}\mapsto f_{\alpha}(|\vec{x}|)$ is the radial function on $\mathbb{R}^{3}$ given by convolving the three-dimensional heat kernel with the harmonic function $\vec{x}\mapsto|\vec{x}|^{-1}$. The equality above follows from the Markov property and the fact that the Bessel process $R$ can be realized as the Euclidean norm of a three-dimensional Brownian motion. A direct computation shows that $\|f_{\alpha}\|_{L^{\infty}(\mathbb{R}^{3})}\leq C(1-\alpha)^{-1/2}$ for a universal constant $C.$
Definition 6.2.
An Airy process of scale $s>0$ centered at $y\in\mathbb{R}$ is the process
$\mathcal{A}_{y}^{s}(x):=s^{1/3}\mathcal{P}_{1}(s^{-2/3}(x-y)),\;\;\;\;\;\;x\in\mathbb{R}$
where $\mathcal{P}_{1}$ is the parabolic Airy2 process; the top curve of the line ensemble in 1.5.
Proposition 6.3.
Fix $\eta\in(0,1)$ and let $\mathcal{A}_{x}^{t}$ and $\tilde{\mathcal{A}_{y}^{s}}$ denote two independent Airy processes of scales $t,s\in(\eta,1]$ centered at $x,y\in\mathbb{R}$ respectively. We also define the notation $\mathcal{V}:=\mathcal{A}_{x}^{t}+\tilde{\mathcal{A}_{y}^{s}}$. Let $I=[a,b]$ be a compact interval. Then $(\mathcal{V}(t+a)-V(a))_{t\in[0,b-a]}$ is absolutely continuous with respect to a Brownian motion of diffusion rate $4$. Furthermore, for all $p\geq 1$, the Radon-Nikodym derivative has a finite $p^{\text{th}}$ moment which is bounded above by $Ce^{Cp^{2}(a^{2}+b^{2})}$ for some universal constant $C=C(p,\eta,x,y)>0$ not depending on $p,a,b,s,t$ such that $0\leq b-a\leq 10$.
The choice $b-a\leq 10$ is arbitrary, but we will never use this bound for intervals larger than length $10$.
Proof.
We reduce the claim to the case where $a=-b.$ Since the Airy2 process is stationary without the addition of the parabola, and since the difference of two quadratic functions is linear, translating the interval $I$ to be centered at the origin amounts to adding a linear drift term of the form $mt+b$ to the process $\mathcal{V}(t)$, for some $m$ whose absolute value can be bounded above by $C(|a|+|b|)$.
The Airy process on $[-a,a]$ with $|a|\leq 10$ is absolutely continuous with respect to a standard Brownian motion on that interval, and moreover the Radon-Nikodym derivative is deterministically bounded due to [Dau23, Corollary 1.3].
Brownian motion with a linear drift of slope $m$ on $[0,b-a]$ is absolutely continuous with respect to Brownian motion with no drift on that same interval, with Radon-Nikodym derivative given by $e^{\pm mB_{b-a}-\frac{m^{2}}{2}(b-a)^{2}}.$ The $p^{\text{th}}$ moment of this random variable is simply $e^{(p^{2}-p)m^{2}(b-a)^{2}}.$ Combining this with the result of the previous paragraph finishes the proof. ∎
Proposition 6.4.
(The crucial bound) Fix $\eta>0$ and recall $\Gamma$ from (4.3). Let $\mathcal{A}_{x}^{t},\tilde{\mathcal{A}_{y}^{s}},\bar{\mathcal{A}_{u}^{a}},\hat{\mathcal{A}_{v}^{b}}$ denote four independent Airy processes of scale $t,s,a,b$ respectively, centered at fixed real numbers $x,y,u,v$ respectively. For any $\delta>0$, there exists a constant $C=C(\eta,x,y,u,v,\delta)>0$ independent of $\delta\in(0,1/3)$, such that we have the following bound uniformly over $\epsilon\in(0,1]$:
$\sup_{\eta\leq s,t,a,b\leq 1}\mathbb{P}\big{(}\Gamma(\mathcal{A}_{x}^{t}+\tilde{\mathcal{A}_{y}^{s}},\bar{\mathcal{A}_{u}^{a}}+\hat{\mathcal{A}_{v}^{b}},\epsilon^{\frac{1}{3}-\delta})\leq\epsilon^{\frac{2}{3}-\delta}\big{)}\leq C\epsilon^{\frac{2}{5}-4\delta}.$
We are not sure whether $\epsilon^{\frac{2}{5}}$ is the optimal bound, however we are unable to improve it through the methods of this paper. The crucial input to proving that the directed landscape is a black noise is the fact that $\frac{2}{5}-\delta$ is larger than $1/3$ for sufficiently small $\delta$. The bound in 6.4 is sufficient because it will give us a bound almost as small as $\epsilon^{2/3}\cdot\epsilon^{2/5}=\epsilon^{16/15}$ in (6.3). This eventually implies a bound in (2.5) with $\varrho$ smaller than $1/15$.
Proof.
In this proof, we leverage the cube-exponential tail decay of the argmaxes of the Airy processes, and combine this with the results of 6.1 and 6.3 to write everything in terms of Bessel processes, where we can apply the bound from 5.4. We make the following definitions:
  • Denote $\mathcal{V}_{1}:=\mathcal{A}_{x}^{t}+\tilde{\mathcal{A}_{y}^{s}}$ and $\mathcal{V}_{2}:=\bar{\mathcal{A}_{u}^{a}}+\hat{\mathcal{A}_{v}^{b}}$ henceforth.
  • Let $G_{1}$ denote the argmax on $\mathbb{R}$ of $\mathcal{V}_{1}$ and let $G_{2}$ denote the argmax on $\mathbb{R}$ of $\mathcal{V}_{2}$.
  • For $k\in\mathbb{Z}$, let $E^{(1)}_{k}$ denote the event that $G_{1}\in[k,k+5]$ and let $E^{(2)}_{k}$ denote the event that $G_{2}\in[k,k+5]$. Note that $E^{(i)}_{k}$ are independent for distinct $i$.
  • For $k_{1},k_{2},\ell\in\mathbb{Z}$ let $F_{k_{1},k_{2},\ell}$ denote the event that $G_{1}\in[k_{1},k_{1}+5]$ , $G_{2}\in[k_{2},k_{2}+5]$, and furthermore the maximum values of both $\mathcal{V}_{1},\mathcal{V}_{2}$ on the interval $[\ell,\ell+5]$ are within 1 of their absolute maximum values on all of $\mathbb{R}.$
  • For $i=1,2$ and $k\in\mathbb{Z}$ let $H^{(i)}_{k}$ denote the event that there exist two points of distance at least $k$ from each other such that the values of $\mathcal{V}_{i}$ at both points are within 1 of $\max_{z\in\mathbb{R}}\mathcal{V}_{i}(z).$ Note that $H^{(i)}_{k}$ are independent for distinct $i$, and
    $F_{k_{1},k_{2},\ell}\subset H^{(1)}_{k_{1}-\ell}\cap H^{(2)}_{k_{2}-\ell}.$
  • For $k\in\mathbb{Z}$ and $i=1,2$, let $G_{i}^{(k)}$ denote the argmax on $[k,k+5]$ of $\mathcal{V}_{i}$ and $M_{k}^{(i)}:=\mathcal{V}_{i}(G^{(k)}_{i})$ denote the maximum value of $\mathcal{V}_{i}$ on $[k,k+5]$.
  • For $\ell\in\mathbb{Z}$ let $K_{\ell}^{\epsilon}$ denote the event that there are points $z_{1}\in[\frac{4}{5}\ell+\frac{1}{5}G_{1}^{(\ell)},\frac{4}{5}(\ell+5)+\frac{1}{5}G_{1}^{(\ell)}]\subset[\ell,\ell+5]$ and $z_{2}\in[\frac{4}{5}\ell+\frac{1}{5}G_{2}^{(\ell)},\frac{4}{5}(\ell+5)+\frac{1}{5}G_{2}^{(\ell)}]\subset[\ell,\ell+5]$ such that $|z_{1}-z_{2}|\leq\epsilon^{\frac{2}{5}}$, and moreover $\mathcal{V}_{i}(z_{i})>M^{(i)}_{\ell}-\epsilon^{\frac{1}{3}-\delta}$ for $i=1,2.$
We use a union bound for the probability that $\Gamma(\mathcal{A}_{x}^{t}+\tilde{\mathcal{A}_{y}^{s}},\bar{\mathcal{A}_{u}^{a}}+\hat{\mathcal{A}_{v}^{b}},\epsilon^{\frac{1}{3}-\delta})\leq\epsilon^{\frac{2}{3}-\delta}.$ The basis for this union bound is the following inclusion:
$\begin{split}\big{\{}\Gamma(\mathcal{A}_{x}^{t}+\tilde{\mathcal{A}_{y}^{s}},\bar{\mathcal{A}_{u}^{a}}+\hat{\mathcal{A}_{v}^{b}},\epsilon^{\frac{1}{3}-\delta})\leq\epsilon^{\frac{2}{3}-\delta}\big{\}}&\subset\bigcup_{k_{1},k_{2},\ell\in\mathbb{Z}}E_{k_{1}}^{(1)}\cap E_{k_{2}}^{(2)}\cap F_{k_{1},k_{2},\ell}\cap K^{\epsilon}_{\ell}\\ &\subset\bigcup_{k_{1},k_{2},\ell\in\mathbb{Z}}E_{k_{1}}^{(1)}\cap E_{k_{2}}^{(2)}\cap H^{(1)}_{k_{1}-\ell}\cap H^{(2)}_{k_{2}-\ell}\cap K^{\epsilon}_{\ell}.\end{split}$ (6.1)
In this equation, we note that $E_{k_{1}}^{(1)}\cap E_{k_{2}}^{(2)}\cap F_{k_{1},k_{2},\ell}=F_{k_{1},k_{2},\ell}$. The purpose of writing the inclusion in this form will be clear in Step 4. The reason for this inclusion is that intervals of the form $I_{\ell}:=\bigcap_{i=1,2}[\frac{4}{5}\ell+\frac{1}{5}G_{i}^{(\ell)},\frac{4}{5}(\ell+4)+\frac{1}{5}G_{i}^{(\ell)}]$ cover all of $\mathbb{R}$ as $\ell$ varies through all of $\mathbb{Z}$, and successive intervals $I_{\ell},I_{\ell+1}$ overlap in an interval of length at least $1$ (this is because each of the two intervals comprising $I_{\ell}$ has length $4$, and therefore $I_{\ell}$ itself has length at least $3$). Consequently, if $z_{1},z_{2}\in\mathbb{R}$ such that $|z_{1}-z_{2}|<\epsilon^{\frac{2}{5}}\leq 1$ then $z_{1},z_{2}\in I_{\ell}$ for some $\ell.$ The union bound yields an infinite sum over $(k_{1},k_{2},\ell)\in\mathbb{Z}^{3}$, and then Hölder’s inequality yields the desired bound. We break this argument into four steps, obtaining individual bounds on the events $E_{k_{1}}^{(1)},E_{k_{2}}^{(2)},H^{(1)}_{k_{1}-\ell},H^{(2)}_{k_{2}-\ell},K^{\epsilon}_{\ell}.$
Step 1. Using the fact that $G_{1},G_{2}$ can be interpreted as points of geodesic locations in the directed landscape, we use 3.1 to obtain $\mathbb{P}(E^{(i)}_{k})\leq Ce^{-d|k|^{3}}$ for some $C,d>0$ independent of $k\in\mathbb{Z}$ and $i=1,2.$
Step 2. We claim that $\mathbb{P}(H^{(i)}_{k})\leq Ce^{-d|k|^{3}}$. This is almost immediate from 4.2, since the parabolic Airy2 process embeds into the Airy sheet by setting $(x,y)=(t,0)$. Either the maximum of $\mathcal{V}_{i}$ is smaller than $1-k$ (which has a probability less than $e^{-c|k|^{3}}$ since the maximum of each $\mathcal{V}_{i}$ is Tracy-Widom GUE distributed with a deterministic shift which depends on $x,y$, which we ignore), or else there exist two points of distance greater than $k$ from each other such that the value of $\mathcal{V}_{1}$ at both points is bigger than $-k$. In order for the latter event to occur, the constant $\mathfrak{C}$ in 4.2 would have to be of order $k^{2}$, which has a probability less than $Ce^{-d|k|^{3}}.$
Step 3. To bound $\mathbb{P}(K_{\ell}^{\epsilon})$, we first note that the event $K_{\ell}^{\epsilon}$ only depends on the behavior of the processes $\mathcal{V}_{i}$ restricted to the interval $[\ell,\ell+5]$, which is the entire point of the union decomposition above. Using 6.3, and using Hölder’s inequality to get rid of the Radon-Nikodym derivative, we obtain $\mathbb{P}(K_{\ell}^{\epsilon})\leq Ce^{Cq^{2}\ell^{2}}\cdot\mathbb{P}_{BM^{\otimes 2}}(K_{\ell}^{\epsilon})^{1/p}$. The latter probability is taken with respect to two independent Brownian motions on $[0,5]$. Under these Brownian motions, the argmaxes $G_{1}^{(\ell)}$ and $G_{2}^{(\ell)}$ have arcsine laws, and we can verify that $\mathbb{P}_{BM^{\otimes 2}}(|G_{1}^{(\ell)}-G_{2}^{(\ell)}|<\epsilon^{\frac{2}{5}})<C\epsilon^{\frac{2}{5}}.$ Therefore,
$\mathbb{P}_{BM^{\otimes 2}}(K_{\ell}^{\epsilon}\cap\{|G_{1}^{(\ell)}-G_{2}^{(\ell)}|<\epsilon^{\frac{2}{5}}\})^{1/p}<\epsilon^{p^{-1}\frac{2}{5}}.$
It only remains to estimate $\mathbb{P}_{BM^{\otimes 2}}(K_{\ell}^{\epsilon}\cap\{|G_{1}^{(\ell)}-G_{2}^{(\ell)}|\geq\epsilon^{\frac{2}{5}}\})$. To do this, we apply the result of 6.1 (with $\alpha=\frac{4}{5},$ since the interval $[\frac{4}{5}\ell+\frac{1}{5}G_{i}^{(\ell)},\frac{4}{5}(\ell+5)+\frac{1}{5}G_{i}^{(\ell)}]\subset[\ell,\ell+5]$ has length $4$) to rewrite the probability in terms of Bessel processes, at the extra cost of a deterministic factor. Rescaling these Bessel processes by $\epsilon^{-2/3}$ in time and by $\epsilon^{1/3}$ in space puts us in the setting of 5.4, with $\gamma=\frac{2}{3}-\frac{2}{5}=\frac{4}{15}.$ We apply that proposition with $\bar{\gamma}=\delta$, where $\delta$ there is chosen the same as $\delta$ here. Applying 5.4 yields a bound of $\epsilon^{\frac{2}{5}-3\delta}$, since $\epsilon^{2/5}=(\epsilon^{4/15})^{3/2}$ and the remaining negative exponents in that proposition can be bounded above by $3\delta$. In summary,
$\mathbb{P}(K_{\ell}^{\epsilon})\leq Ce^{C\ell^{2}}\epsilon^{p^{-1}(\frac{2}{5}-3\delta)},$
where $C$ is uniform over $\ell,\epsilon$.
Step 4. Finally, we apply the union bound in (6.1), then use Hölder’s inequality with some other conjugate exponents $p^{\prime},q^{\prime}$ (possibly different from the exponents $p,q$ from earlier) and thanks to the results of Steps 1-3 we obtain
$\mathbb{P}\big{(}\Gamma(\mathcal{A}_{x}^{t}+\tilde{\mathcal{A}_{y}^{s}},\bar{\mathcal{A}_{u}^{a}}+\hat{\mathcal{A}_{v}^{b}},\epsilon^{\frac{1}{3}-\delta})\leq\epsilon^{\frac{2}{3}-\delta}\big{)}$ $\leq\sum_{k_{1},k_{2},\ell\in\mathbb{Z}}\mathbb{P}(E_{k_{1}}^{(1)}\cap E_{k_{2}}^{(2)}\cap H^{(1)}_{k_{1}-\ell}\cap H^{(2)}_{k_{2}-\ell}\cap K^{\epsilon}_{\ell})$
$\leq\sum_{k_{1},k_{2},\ell\in\mathbb{Z}}\mathbb{P}(E_{k_{1}}^{(1)}\cap E_{k_{2}}^{(2)})^{\frac{1}{2q^{\prime}}}\mathbb{P}(H^{(1)}_{k_{1}-\ell}\cap H^{(2)}_{k_{2}-\ell})^{\frac{1}{2q^{\prime}}}\mathbb{P}(K^{\epsilon}_{\ell})^{\frac{1}{p^{\prime}}}$
$\leq\epsilon^{\frac{1}{pp^{\prime}}(\frac{2}{5}-3\delta)}\sum_{k_{1},k_{2},\ell\in\mathbb{Z}}e^{-\frac{c}{2q^{\prime}}\left(|k_{1}|^{3}+|k_{2}|^{3}+|\ell-k_{1}|^{3}+|\ell-k_{2}|^{3}\right)}\cdot e^{C\frac{q^{2}}{p^{\prime}}\ell^{2}}.$
For the term $e^{C\frac{q^{2}}{p^{\prime}}\ell^{2}},$ we use the bound $\ell^{2}=\frac{1}{2}\ell^{2}+\frac{1}{2}\ell^{2}\leq k_{1}^{2}+(\ell-k_{1})^{2}+k_{2}^{2}+(\ell-k_{2})^{2}.$ We perform the sums in the expression above, first over $\ell$ and then over $(k_{1},k_{2})$. The cube-exponential decay overwhelms the square-exponential growth terms in each case, which shows that the infinite sum converges to a finite value for any (fixed) choice of $p,p^{\prime}$. We conclude the proof by taking $p,p^{\prime}$ very close to $1$ (and thus $q,q^{\prime}$ large) so that $\frac{1}{pp^{\prime}}(\frac{2}{5}-3\delta)=\frac{2}{5}-4\delta.$
Lemma 6.5.
Fix $\eta\in(0,1)$. Let $\mathcal{A}_{x}^{t}$ and $\tilde{\mathcal{A}_{y}^{s}}$ denote two independent Airy processes of scale $t,s$ centered at $x,y\in\mathbb{R}$ respectively, and denote $\mathcal{V}:=\mathcal{A}_{x}^{t}+\tilde{\mathcal{A}_{y}^{s}}$. As in Eq. 4.2, let $N_{\epsilon}(\mathcal{V})$ denote the smallest number of intervals of length $\epsilon^{2}$ needed to cover the random set $\mathscr{S}(\mathcal{V},\epsilon).$ Then uniformly over all scales $s,t\in[\eta,1]$, $\sup_{\epsilon\in(0,1]}\mathbb{E}[N_{\epsilon}(\mathcal{V})^{p}]<\infty$ for all $p\geq 1.$
Proof.
Let $N^{k}_{\epsilon}(\mathcal{V})$ denote the smallest number of intervals contained in $[k,k+2]$ that are needed to cover the random set $\mathscr{S}(\mathcal{V},\epsilon)\cap[k,k+2].$ We have a trivial bound $N_{\epsilon}(\mathcal{V})\leq\sum_{k\in\mathbb{Z}}N^{k}_{\epsilon}(\mathcal{V})$ and Minkowski’s inequality implies
$\sup_{\epsilon\in(0,1]}\mathbb{E}[N_{\epsilon}(\mathcal{V})^{p}]^{1/p}\leq\sum_{k\in\mathbb{Z}}\sup_{\epsilon\in(0,1]}\mathbb{E}[N^{k}_{\epsilon}(\mathcal{V})^{p}]^{1/p}.$ (6.2)
Let $E_{k}$ denote the event that $N^{k}_{\epsilon}(\mathcal{V})$ is nonzero, that is, the event that $\mathscr{S}(\mathcal{V},\epsilon)\cap[k,k+2]$ is nonempty. Then either the maximum of $\mathcal{V}$ on all $\mathbb{R}$ is smaller than $1-k$ (which has a probability less than $e^{-d|k|^{3}}$ since we know that the maximum of $\mathcal{V}$ is Tracy-Widom GUE distributed plus a deterministic shift depending only on $x,y$ which we are ignoring) or else there exists a point in $[k,k+2]$ such that the value of $\mathcal{V}$ at this point is bigger than $-k$ (since $\epsilon\leq 1$ by assumption). But the constant $\mathfrak{C}$ in 4.2 would have to be of order $k^{2}$ for the latter event to occur, which has a probability less than $Ce^{-d|k|^{3}}.$ Thus $\mathbb{P}(E_{k})\leq Ce^{-d|k|^{3}}$ for some $C,d>0$.
We can bound $N_{\epsilon}^{k}(\mathcal{V})$ by the number $M_{\epsilon}^{k}(\mathcal{V})$ of intervals of size $\epsilon^{2}$ contained in $[k-1,k+3]$ needed to cover the set of points $z\in[k,k+2]$ such that $\mathcal{V}(z)$ is within $\epsilon$ of its maximum value on $[k-1,k+3]$. We apply 6.3 and 6.1 with $\alpha=3/4,$ and see that up to a deterministic factor of size $Ce^{Ck^{2}}$ and possibly enlarging $p$ due to Hölder’s inequality, the process $\mathcal{V}$ can be replaced by a standard two-sided Bessel process. By this observation and 5.2 we conclude that $\mathbb{E}[M_{\epsilon}^{k}(\mathcal{V})^{p}]\leq Ce^{Ck^{2}}$ for some absolute constants $C,d$ independent of $k,\epsilon$ but possibly depending on $p$. We find that
$\mathbb{E}[N^{k}_{\epsilon}(\mathcal{V})^{p}]=\mathbb{E}[N^{k}_{\epsilon}(\mathcal{V})^{p}1_{E_{k}}]\leq\mathbb{E}[M^{k}_{\epsilon}(\mathcal{V})^{2p}]^{1/2p}\mathbb{P}(1_{E_{k}})^{1/2}\leq Ce^{-d|k|^{3}}\cdot Ce^{Ck^{2}},$
where $C,d$ are uniform over all $k,\epsilon.$ We sum over $k\in\mathbb{Z}$ and use (6.2) to obtain the desired bound. ∎
Lemma 6.6.
Let $\mathcal{S}$ be the Airy sheet, and for a function $f:\mathbb{R}^{2}\to\mathbb{R}$ let $[f]_{\alpha,M}:=\sup_{\vec{s}\neq\vec{t}\in[-M,M]^{2}}|\vec{t}-\vec{s}|^{-\alpha}|f(\vec{t})-f(\vec{s})|$ denote the $\alpha$-Hölder semi-norm of $f$ on $[-M,M]^{2}$. For any $\alpha<1/2$ and $\delta\in(0,1/2)$, there exists $C,d>0$ such that uniformly over $M>0$ and $u>0$,
$\mathbb{P}([\mathcal{S}]_{\alpha,M}>u)\leq CM^{10}e^{-dM^{-\delta}u^{3/2}}.$
Proof.
This follows from [DOV22, Proposition 10.5], by taking $\tau=0$ and observing that $\log^{1/2}(b\xi^{-1})\leq C(\delta)b^{\delta}\xi^{-\delta}$ for arbitrary $\delta,\xi>0$. ∎
Proposition 6.7 (Controlling the error bounds).
We fix $\eta>0$ and recall the error terms $\mathfrak{E}_{\epsilon}(f,B)$ from 4.9. We consider two independent Airy processes $\mathcal{A}^{s}_{x},\mathcal{A}^{t}_{y}$ of scales $s,t\in[\eta,1]$ and centered at $x,y\in\mathbb{R}$ respectively, and let $\mathcal{V}:=\mathcal{A}^{s}_{x}+\mathcal{A}^{t}_{y}$. We define the random variable
$\mathfrak{B}:=\inf\{B>0:-B-2\eta^{-1}u^{2}\leq\mathcal{V}(u)\leq B-2\eta u^{2},\text{ for all }u\in\mathbb{R}\}.$
Fix $p\geq 1$. Then $\mathbb{E}[\mathfrak{B}^{p}]<\infty$, and there exist $C,d>0$ depending only on $\eta,p$ (and not on $s,t,\epsilon$) such that
$\mathbb{E}[\mathfrak{E}_{\epsilon}(\mathcal{V},\mathfrak{B})^{p}]\leq Ce^{-d\log^{6/5}(1/\epsilon)}.$
Proof.
By 4.2, the random variable $\mathfrak{B}$ is almost surely finite, and in fact satisfies $\mathbb{P}(\mathfrak{B}>u)\leq Ce^{-du^{3/2}}$ for some $C,d>0$ that may depend on $\eta$ and $x_{i},y_{i}$ (but not on $s,t$). We see that $\mathbb{E}[1_{\{\mathfrak{B}>d\epsilon^{-1/3}\}}]\leq Ce^{-d\epsilon^{-1}}$ where the constants $C,d$ may have decreased or increased (respectively).
It remains to bound the Hölder semi-norm term in the expression for $\mathfrak{E}_{\epsilon}$. We set $\tilde{\delta}:=\frac{3}{40}\delta$, and note that the probability of $100\mathfrak{B}>\log^{4/5}(1/\epsilon)$ decays like $Ce^{-d\log^{6/5}(1/\epsilon)}$. Consequently, we just need to show that
$\mathbb{P}\bigg{(}\mathfrak{Hol}\big{(}\mathcal{V},\frac{1}{2}(1-\tilde{\delta}),1+2\log^{4/5}(1/\epsilon)\big{)}>\epsilon^{-\tilde{\delta}/3}\bigg{)}\leq Ce^{-d\log^{6/5}(1/\epsilon)}.$
We use 6.6 to see that the left side is bounded above by $C\log^{40/5}(1/\epsilon)\cdot e^{-d\log^{-4\delta/5}(1/\epsilon)\cdot\epsilon^{-\tilde{\delta}/2}}.$ This is far smaller than the required bound. ∎
Now we are finally in a position to prove (2.5).
Theorem 6.8.
The estimate (2.5) holds for any $\varrho<\frac{1}{15}$, consequently the directed landscape is a black noise.
Proof.
Using the metric composition property of the directed landscape, we can write the conditional expectation inside the variance of (2.5) for $0\leq a<b\leq 1$ as
$\mathbb{E}\left[\prod_{i=1}^{n}\mathcal{L}_{0,1}(x_{j},y_{j})\bigg{|}\mathcal{F}^{\mathcal{L}}_{a,b}\right]$ $=\mathbb{E}\left[\prod_{i=1}^{n}f_{i}^{a}\circ\mathcal{L}_{a,b}\circ g_{i}^{b}\bigg{|}\mathcal{F}_{a,b}^{\mathcal{L}}\right]$
$=\int_{C(\mathbb{R})^{n}}\int_{C(\mathbb{R})^{n}}\prod_{i=1}^{n}(u_{i}\circ\mathcal{L}_{a,b}\circ v_{i})\nu_{a}(d\vec{u})\mu_{b}(d\vec{v}),$
where $f_{i}^{t}:=f_{i}\circ_{\ell}\mathcal{L}_{0,t}$ and $g_{i}^{t}:=\mathcal{L}_{t,1}\circ_{r}g_{i},$ where the left and right metric compositions are defined as
$(f\circ_{\ell}L)(y):=\sup_{x\in\mathbb{R}}f(x)+L(x,y),$ $(L\circ_{r}g)(x)=\sup_{y\in\mathbb{R}}L(x,y)+g(y),$
and where $\nu_{a}$ and $\mu_{b}$ are respectively the laws of $(f_{1}^{a},...,f_{n}^{a})$ and $(g^{b}_{1},...,g^{b}_{n}).$ Fix $\eta,\delta\in(0,1/2)$. From the last expression we deduce that to prove the theorem, it suffices to show that
$\int_{(C(\mathbb{R})^{n})^{4}}\mathrm{Cov}\left(\prod_{i=1}^{n}(f_{i}\circ\mathcal{L}_{0,b-a}\circ g_{i}),\prod_{j=1}^{n}(u_{j}\circ\mathcal{L}_{0,b-a}\circ v_{j})\right)\nu_{a}^{\otimes 2}(d\vec{f},d\vec{u})\mu_{b}^{\otimes 2}(d\vec{g},d\vec{v})\leq C(b-a)^{\frac{16}{15}-\delta}.$ (6.3)
where $\eta\leq a<b\leq 1-\eta$ and $C$ is allowed to depend upon $\eta,\delta\in(0,1/2)$ and on $x_{j},y_{j}$ but not on $a,b$.
In this proof, we abbreviate $[n]:=\{1,...,n\}$. By taking the Taylor expansion of the function $(x_{1},...,x_{n})\mapsto x_{1}\cdots x_{n}$ around some fixed value $(a_{1},...,a_{n}),$ we rewrite it as $\sum_{S\subset[n]}a_{S}(x-a)_{S^{c}},$ where $x_{S}=x_{i_{1}}\cdots x_{i_{r}}$ if $S=\{i_{1},...,i_{r}\}.$ We apply this fact with $x_{i}:=f_{i}\circ\mathcal{L}_{0,\epsilon}\circ g_{i},$ and $a_{i}:=\max(f_{i}+g_{i}),$ which allows us to rewrite the products in (6.3) as linear combinations of quantities of the form $\max(f+g)_{S}\cdot(f\circ\mathcal{L}_{0,\epsilon}\circ g-\max(f+g))_{S^{c}},$ whereas before $S$ is a subset of $[n]$ and the subscript denotes the product of the respective quantity over all indices which lie in $S$. Consequently, to prove (6.3) it suffices to show that for all subsets $S,T\subset[n]$,
$\begin{split}\int_{(C(\mathbb{R})^{n})^{4}}\mathrm{Cov}\bigg{(}(f\circ\mathcal{L}_{0,\epsilon}\circ g&-\max(f+g))_{S^{c}},(u\circ\mathcal{L}_{0,\epsilon}\circ v-\max(u+v))_{T^{c}}\bigg{)}\\ &\cdot\max(f+g)_{S}\max(u+v)_{T}\nu_{a}^{\otimes 2}(d\vec{f},d\vec{u})\mu_{b}^{\otimes 2}(d\vec{g},d\vec{v})\leq C\epsilon^{\frac{16}{15}-\delta}.\end{split}$ (6.4)
Under $\nu_{a}^{\otimes 2}\otimes\mu_{b}^{\otimes 2}$ the $f_{i},g_{i},u_{i},v_{i}$ are all distributed as independent parabolic Airy2 processes of different scales (these scales are always in $[\eta,1-\eta]$), perhaps recentered at some values depending on $x_{i},y_{i}$ (which are fixed). Consequently, the individual maxima of $f_{i}+g_{i}$ and $u_{i}+v_{i}$ each have Tracy-Widom GUE laws (perhaps rescaled and recentered). Therefore,
$\sup_{\eta\leq a<b\leq 1-\delta}\max_{S,T\subset[n]}\int_{(C(\mathbb{R})^{n})^{4}}\bigg{|}\max(f+g)_{S}\max(u+v)_{T}\bigg{|}^{q}\nu_{a}^{\otimes 2}(d\vec{f},d\vec{u})\mu_{b}^{\otimes 2}(d\vec{g},d\vec{v})<\infty.$
By Hölder’s inequality (choosing $p$ very close to 1 and thus $q$ large), to prove (6.4) it suffices to show that for all $S,T\subset[n]$ we have
$\bigg{[}\int_{(C(\mathbb{R})^{n})^{4}}\bigg{|}\mathrm{Cov}\bigg{(}(f\circ\mathcal{L}_{0,\epsilon}$ $\circ g-\max(f+g))_{S^{c}},$
$(u\circ\mathcal{L}_{0,\epsilon}\circ v-\max(u+v))_{T^{c}}\bigg{)}\bigg{|}^{p}\nu_{a}^{\otimes 2}(d\vec{f},d\vec{u})\mu_{b}^{\otimes 2}(d\vec{g},d\vec{v})\bigg{]}^{1/p}\leq C\epsilon^{\frac{16}{15}-\delta}.$ (6.5)
Let #$\#S$ denote the cardinality of the set $S$. If #$\#(S^{c})+\#(T^{c})\geq 4,$ then we may apply 4.11 to obtain a bound of size $\epsilon^{\frac{4}{3}-\delta}$ plus some number of “error terms” which by 6.7 have a super-polynomial rate of decay as $\epsilon\to 0$. This achieves the goal in (6.5).
By a symmetry argument, what remains is to prove (6.5) in the case when (#$(\#(S^{c}),\#(T^{c}))=(1,1)$ or (#$(\#(S^{c}),\#(T^{c}))=(1,2)$. We consider the case when (#$(\#(S^{c}),\#(T^{c}))=(1,1)$. In this case, the covariance in the integrand reduces to $\operatorname{Cov}(f_{1}\circ\mathcal{L}_{0,\epsilon}\circ g_{1},u_{1}\circ\mathcal{L}_{0,\epsilon}\circ v_{1})$. We apply 4.9 and 4.11 (with $k+\ell=2$) and we obtain (modulo some “error terms” which by 6.7 have a super-polynomial rate of decay as $\epsilon\to 0$) an upper bound given by
$\bigg{[}\int_{(C(\mathbb{R})^{n})^{4}}$ $\mathfrak{B}^{2\delta}\min\bigg{\{}\epsilon^{\frac{2}{3}-\delta},$
$\big{(}N_{\epsilon^{\frac{1}{3}-\delta}}(f_{1}+g_{1})+N_{\epsilon^{\frac{1}{3}-\delta}}(u_{1}+v_{1})\big{)}\cdot e^{-c\epsilon^{-2}\Gamma\big{(}f_{1}+g_{1},u_{1}+v_{1};\epsilon^{\frac{1}{3}-\delta}\big{)}^{3}}\bigg{\}}\nu_{a}^{\otimes 2}(d\vec{f},d\vec{u})\mu_{b}^{\otimes 2}(d\vec{g},d\vec{v})\bigg{]}^{1/p}.$
By 6.7 and another application of Hölder’s inequality with $p$ very close to $1$, we can ignore the factor of $\mathfrak{B}^{2\delta}$, since $\mathbb{E}[\mathfrak{B}^{2\delta q}]<\infty$ for the conjugate exponent $q$. We use the elementary bound $\min\{a,b\}\leq a^{\theta}b^{1-\theta}$ for $\theta\in[0,1]$, and see that the last expression is bounded above by
$\bigg{[}\int_{(C(\mathbb{R})^{n})^{4}}\epsilon^{(1-\theta)(\frac{2}{3}-\delta)}$ $\cdot\big{(}N_{\epsilon^{\frac{1}{3}-\delta}}(f_{1}+g_{1})+N_{\epsilon^{\frac{1}{3}-\delta}}(u_{1}+v_{1})\big{)}^{\theta}$
$\cdot e^{-c\theta\epsilon^{-2}\Gamma\big{(}f_{1}+g_{1},u_{1}+v_{1};\epsilon^{\frac{1}{3}-\delta}\big{)}^{3}}\nu_{a}^{\otimes 2}(d\vec{f},d\vec{u})\mu_{b}^{\otimes 2}(d\vec{g},d\vec{v})\bigg{]}^{1/p},$ (6.6)
for all $\theta\in[0,1]$. We split the integral into two parts according to $\Gamma(f_{1}+g_{1},u_{1}+v_{1},\epsilon^{\frac{1}{3}-\delta})<\epsilon^{\frac{2}{3}-\delta}$ or $\Gamma(u_{1}+v_{1},f_{1}+g_{1},\epsilon^{\frac{1}{3}-\delta})\geq\epsilon^{\frac{2}{3}-\delta}.$ On the latter event we obtain an exponential rate of decay of order $e^{-c\epsilon^{-3\delta}},$ which decays to 00 super-polynomially fast as $\epsilon\to 0.$ In the case that $\Gamma(f_{1}+g_{1},u_{1}+v_{1},\epsilon^{\frac{1}{3}-\delta})<\epsilon^{\frac{2}{3}-\delta}$, we apply 6.5 and Hölder’s inequality (with $p$ close to 1 and $q$ large) to argue that we can disregard the term $\big{(}N_{\epsilon^{\frac{1}{3}-\delta}}(f_{1}+g_{1})+N_{\epsilon^{\frac{1}{3}-\delta}}(u_{1}+v_{1})\big{)}^{\theta}$ at the cost of making the value of $p$ slightly larger. Then we apply 6.4 to see that the integral in (6.6) will decay as $\epsilon^{(1-\theta)(\frac{2}{3}-\delta)}\cdot\epsilon^{\frac{2}{5}-4\delta}$. In this setting, we can make the exponent on $\epsilon$ as close to $\frac{16}{15}$ as desired by taking $\theta,\delta$ close to 00 and $p$ close to $1$.
Next, we consider the case when (#$(\#(S^{c}),\#(T^{c}))=(1,2)$. The argument in this case is very similar, we apply 4.10 instead of 4.9, and we still apply 4.11, in this case with $k+\ell=3$. In (6.6) this gives (modulo the error terms) a bound of size
$\bigg{[}\int_{(C(\mathbb{R})^{n})^{4}}\epsilon^{(1-\theta)(1-\delta)}\cdot\big{(}N_{\epsilon^{\frac{1}{3}-\delta}}(f_{1}+g_{1})+N_{\epsilon^{\frac{1}{3}-\delta}}(u_{1}$ $+v_{1})+N_{\epsilon^{\frac{1}{3}-\delta}}(u_{2}+v_{2})\big{)}^{\theta}$
$\cdot e^{-c\theta\epsilon^{-2}\Gamma(\epsilon)^{3}}\nu_{a}^{\otimes 2}(d\vec{f},d\vec{u})\mu_{b}^{\otimes 2}(d\vec{g},d\vec{v})\bigg{]}^{1/p},$ (6.7)
where $\Gamma(\epsilon):=\min\{\Gamma\big{(}f_{1}+g_{1},u_{1}+v_{1};\epsilon^{\frac{1}{3}-\delta}\big{)},\Gamma\big{(}f_{1}+g_{1},u_{2}+v_{2};\epsilon^{\frac{1}{3}-\delta}\big{)}\}.$ We split the integral into two parts according to $\Gamma(\epsilon)<\epsilon^{\frac{2}{3}-\delta}$ or $\Gamma(\epsilon)\geq\epsilon^{\frac{2}{3}-\delta}.$ The latter of these terms yields a super-polynomially fast rate of decay to zero in (6.7), whereas the former can be bounded by 5.2 and 6.4, noting from 6.4 and a union bound that
$\mathbb{P}(\Gamma(\epsilon)<\epsilon^{\frac{2}{3}-\delta})\leq\mathbb{P}\big{(}\Gamma\big{(}f_{1}+g_{1},u_{1}+v_{1};\epsilon^{\frac{1}{3}-\delta}\big{)}<\epsilon^{\frac{2}{3}-\delta}\big{)}+\mathbb{P}\big{(}\Gamma\big{(}f_{1}+g_{1},u_{2}+v_{2};\epsilon^{\frac{1}{3}-\delta}\big{)}<\epsilon^{\frac{2}{3}-\delta}\big{)}\leq C\epsilon^{\frac{2}{5}-4\delta}.$
This will yield an expression at most $\epsilon^{(1-\theta)(1-\delta)}\cdot\epsilon^{\frac{2}{5}-4\delta}$ in (6.7). By choosing $\delta,\theta$ close to 00, the exponent on $\epsilon$ can be made as close to $\frac{7}{5}$ as desired by taking $\theta$ close to 00, which exceeds the desired threshold of $\epsilon^{\frac{16}{15}-\delta}$. ∎
7. Independence from Gaussian white noise
In this section, we prove 1.11.
Proof of 1.11.
Assume that we have a probability space equipped with a noise, $(\Omega,(\mathcal{F}_{s,t})_{s<t},P,\theta)$, and that $\mathcal{L},\xi$ are (respectively) a directed landscape and a standard white noise which are adapted to (and together generate) the filtration $\mathcal{F}_{s,t}$,
$\mathcal{F}_{s,t}:=\mathcal{F}^{\xi}_{s,t}\vee\mathcal{F}^{\mathcal{L}}_{s,t},$
where $\mathcal{F}_{s,t}^{\xi}$ is the filtration generated by the white noise and $\mathcal{F}_{s,t}^{\mathcal{L}}$ is the filtration generated by the directed landscape. Our goal is to show that $\mathcal{F}^{\xi},\mathcal{F}^{\mathcal{L}}$ are independent of each other. We will break the proof down into five steps.
Step 1. For $h\in L^{2}(\mathbb{R}^{2})$, we define the pairing $(h,\xi):=\int_{\mathbb{R}^{2}}h(t,x)\xi(dt,dx).$ We claim that these are linear random variables on this probability space, in fact we have that $E[(h,\xi)|\mathcal{F}_{s,t}]=(h1_{[s,t]\times\mathbb{R}},\xi)$. This would be obvious if $\mathcal{F}^{\xi}_{s,t}$ was the full $\sigma$-algebra $\mathcal{F}_{s,t}$, but needs some verification in the present context, where $\mathcal{F}_{s,t}$ may contain strictly more information than $\mathcal{F}^{\xi}_{s,t}$.
Note that $(h1_{[s,t]\times\mathbb{R}},\xi)$ is $\mathcal{F}_{s,t}$-measurable by assumption, therefore the definition of a noise (1.3) implies that it is independent of $\mathcal{F}_{t,u}$ and $\mathcal{F}_{a,s}$ for any $a<s<t<u$. From this fact, we can prove the claim by writing the almost sure identity
$(h,\xi)=(h1_{(-\infty,s]\times\mathbb{R}},\xi)+(h1_{[s,t]\times\mathbb{R}},\xi)+(h1_{[t,\infty)\times\mathbb{R}]},\xi),$
and then noting that the three terms on the right are respectively $\mathcal{F}_{-\infty,s}$-measurable, $\mathcal{F}_{s,t}$-measurable, and $\mathcal{F}_{t,\infty}$-measurable. By taking the conditional expectation given $\mathcal{F}_{s,t}$, and using independence of these three $\sigma$-algebras, the first and third term vanish while the second term is unaffected by the conditional expectation.
Step 2. We argue that $(h,\xi)^{\mathcal{L}}:=E[(h,\xi)|\mathcal{F}^{\mathcal{L}}]$ is a linear random variable on the restricted noise $(\Omega,(\mathcal{F}_{s,t}^{\mathcal{L}})_{s<t},P,\theta)$ for any $h\in L^{2}(\mathbb{R}^{2})$. The black noise property of $\mathcal{L}$ proved in 1.10 will then imply that $E[(h,\xi)|\mathcal{F}^{\mathcal{L}}]=0$ for any $h\in L^{2}(\mathbb{R}^{2})$. For $s<t<u$ we have
$\mathbb{E}\big{[}(h,\xi)^{\mathcal{L}}|\mathcal{F}^{\mathcal{L}}_{s,u}\big{]}$ $=\mathbb{E}\big{[}(h,\xi)|\mathcal{F}^{\mathcal{L}}_{s,u}\big{]}$
$=\mathbb{E}\big{[}\mathbb{E}[(h,\xi)|\mathcal{F}_{s,u}]|\mathcal{F}^{\mathcal{L}}_{s,u}\big{]}$
$=\mathbb{E}\bigg{[}\big{(}\mathbb{E}[(h,\xi)|\mathcal{F}_{s,t}]+\mathbb{E}[(h,\xi)|\mathcal{F}_{t,u}]\big{)}\bigg{|}\mathcal{F}^{\mathcal{L}}_{s,u}\bigg{]}$
$=\mathbb{E}\big{[}\mathbb{E}[(h,\xi)|\mathcal{F}_{s,t}]\big{|}\mathcal{F}^{\mathcal{L}}_{s,u}\big{]}+\mathbb{E}\big{[}\mathbb{E}[(h,\xi)|\mathcal{F}_{t,u}]\big{|}\mathcal{F}^{\mathcal{L}}_{s,u}\big{]}$
$=\mathbb{E}\big{[}\mathbb{E}[(h,\xi)|\mathcal{F}_{s,t}]\big{|}\mathcal{F}^{\mathcal{L}}_{s,t}\big{]}+\mathbb{E}\big{[}\mathbb{E}[(h,\xi)|\mathcal{F}_{t,u}]\big{|}\mathcal{F}^{\mathcal{L}}_{t,u}\big{]}$
$=\mathbb{E}\big{[}(h,\xi)\big{|}\mathcal{F}^{\mathcal{L}}_{s,t}\big{]}+\mathbb{E}\big{[}(h,\xi)\big{|}\mathcal{F}^{\mathcal{L}}_{t,u}\big{]}$
$=\mathbb{E}\big{[}(h,\xi)^{\mathcal{L}}|\mathcal{F}^{\mathcal{L}}_{s,t}\big{]}+\mathbb{E}\big{[}(h,\xi)^{\mathcal{L}}|\mathcal{F}^{\mathcal{L}}_{t,u}\big{]}.$
In the first and second line, we simply used the tower property of conditional expectation. In the third line, we used the result of Step 1. In the fifth line, we are using the fact that $\mathcal{F}^{\mathcal{L}}_{s,t},\mathbb{E}[(h,\xi)|\mathcal{F}_{s,t}]$ are independent of $\mathcal{F}^{\mathcal{L}}_{t,u}\subset\mathcal{F}_{t,u}$, and likewise we are using the fact that $\mathcal{F}^{\mathcal{L}}_{t,u},\mathbb{E}[(h,\xi)|\mathcal{F}_{t,u}]$ are independent of $\mathcal{F}_{s,t}^{\mathcal{L}}\subset\mathcal{F}_{s,t}$. In the sixth and seventh line, we simply used the tower property of conditional expectation.
Step 3. Let $h_{s_{j},t_{j}}\in L^{2}(\mathbb{R}^{2})$ be supported on $(s_{j},t_{j})\times\mathbb{R}$, where $(s_{j},t_{j})$ are disjoint intervals. Using Step 2 and the noise property,
$\mathbb{E}\bigg{[}\prod_{j=1}^{n}F_{s_{j},t_{j}}(\mathcal{L})\cdot(h_{s_{j},t_{j}},\xi)\bigg{]}=\prod_{j=1}^{n}\mathbb{E}\bigg{[}F_{s_{j},t_{j}}(\mathcal{L})\cdot(h_{s_{j},t_{j}},\xi)\bigg{]}=0,$
where $F_{s_{j},t_{j}}:C(\mathbb{R}^{4}_{\uparrow})\to\mathbb{R}$ are arbitrary bounded functionals which are $\mathcal{F}_{s_{j},t_{j}}^{\mathcal{L}}$-measurable for $1\leq j\leq n$, and we use the result of Step 2 in the last equality.
Step 4. As before, let $h_{s_{j},t_{j}}\in L^{2}(\mathbb{R}^{2})$ be supported on $(s_{j},t_{j})\times\mathbb{R}$, where $(s_{j},t_{j})$ are disjoint intervals. Using standard stochastic calculus tools for the Brownian motion, we will show that linear combinations of random variables of the form $c+\prod_{j=1}^{n}(h_{s_{j},t_{j}},\xi)$ are actually dense in $L^{2}(\Omega,\mathcal{F}^{\xi},P)$, as we vary over all $c\in\mathbb{R}$, $n\in\mathbb{N}$ and disjoint intervals $(s_{j},t_{j})$.
There are many ways to prove the desired claim, but we will take the following approach. First choose an orthonormal basis $\{e_{j}\}_{j\geq 1}$ of $L^{2}(\mathbb{R})$ and define the independent, identically distributed Brownian motions $B^{j}(t):=(1_{[0,t]}\otimes e_{j},\xi).$ Note that $\mathcal{F}^{\xi}_{s,t}$ is precisely the $\sigma$-algebra generated by $(B^{j}(u)-B^{j}(s))_{u\in[s,t],j\in\mathbb{N}}.$ Fix some natural numbers $\ell$ and $p_{1}<...<p_{\ell}$. Let $t^{n}_{j}=j2^{-n}$, and notice that by standard construction of stochastic integrals we have (as $n\to\infty$) the convergence
$\sum_{\sigma\in S_{p_{\ell}}}\sum_{\begin{subarray}{c}s2^{n}\leq k_{1}<k_{2}<...<k_{p_{\ell}}\leq t2^{n}\\ k_{i}\in\mathbb{Z}\end{subarray}}\prod_{j=1}^{\ell}\prod_{r=p_{j-1}}^{p_{j}-1}\big{(}B^{j}(t^{n}_{k_{\sigma(r)}+1})-B^{j}(t^{n}_{k_{\sigma(r)}})\big{)}$
$\stackrel{{\scriptstyle L^{2}(\Omega)}}{{\longrightarrow}}\sum_{\sigma\in S_{p_{\ell}}}$ $\int_{s\leq s_{1}<s_{2}<...<s_{p_{\ell}}\leq t}\prod_{j=1}^{\ell}\prod_{r=p_{j-1}}^{p_{j}-1}dB^{j}(s_{k_{\sigma(r)}+1})=\prod_{j=1}^{\ell}H_{p_{j}-p_{j-1}}\big{(}(t-s)^{-1/2}(B^{j}(t)-B^{j}(s))\big{)},$
where $H_{p}(x)$ is the $p^{th}$ standard Hermite polynomial and $S_{k}$ is the symmetric group of permutations of $\{1,...,k\}$. But for fixed $s<t$, the random variables of the form of the right side form an orthonormal set whose closed linear span in $L^{2}(\Omega)$ (as we vary over natural numbers $\ell$ and $0=p_{0}<...<p_{\ell},$) includes all random variables of the form $\varphi\big{(}B^{1}(t)-B^{1}(s),...,B^{n}(t)-B^{n}(s)\big{)}$ where $n\in\mathbb{N}$ and $\varphi:\mathbb{R}^{n}\to\mathbb{R}$ is bounded measurable, by e.g. the standard theory of Gaussian chaos [Nua06]. Consequently, the linear span $E$ in $L^{2}(\Omega)$ of random variables of the form $c+\prod_{j=1}^{n}(h_{s_{j},t_{j}},\xi)$ contains any random variable of the form $\prod_{j=1}^{m}\varphi_{j}\big{(}B^{1}(t_{j})-B^{1}(s_{j}),...,B^{n}(t_{j})-B^{n}(s_{j})\big{)}$, where $m,n\in\mathbb{N}$ and $\varphi_{j}:\mathbb{R}^{n}\to\mathbb{R}$ are bounded measurable and $(s_{j},t_{j})$ are disjoint intervals. Therefore, by the tensorization property of $L^{2}$ spaces (explained in the proof of 2.4) the linear span $E$ contains all random variables of the form $\Psi\big{(}(B^{i}(t_{j})-B^{i}(s_{j}))_{1\leq j\leq m,1\leq i\leq n}\big{)}$, where $m,n\in\mathbb{N}$ and $\Psi:\mathbb{R}^{m\times n}\to\mathbb{R}$ is bounded measurable and $(s_{j},t_{j})$ are disjoint intervals with dyadic endpoints. This proves the claim, as random variables of the latter form can be used to approximate arbitrary functionals $G(\xi)$, by letting $m,n\uparrow\infty$ and using e.g. the result of 2.1.
Step 5. Using the results of Steps 3 and 4, we apply a density argument to show that $E[F(\mathcal{L})G(\xi)]=E[F(\mathcal{L})]E[G(\xi)]$ for all bounded measurable $F,G$ on the respective canonical spaces, completing the proof. By the result of Step 4, it suffices to prove the claim when $G(\xi)=c+\prod_{j=1}^{r}(h_{s_{j},t_{j}},\xi),$ where $c\in\mathbb{R}$ and $(s_{j},t_{j})$ are disjoint intervals with endpoints taking values in $\{k2^{-n}:k\in\mathbb{Z}\}$ for some $n\in\mathbb{N}.$
In turn, it suffices to prove the claim when $F$ is a finite linear combination of functionals of the form $\prod_{j=0}^{r+1}F_{s_{j},t_{j}}(\mathcal{L})$ where $F_{s_{j},t_{j}}:C(\mathbb{R}^{4}_{\uparrow})\to\mathbb{R}$ are $\mathcal{F}_{s_{j},t_{j}}^{\mathcal{L}}$-measurable and the extreme endpoints are defined by $s_{0}:=-\infty$, $t_{0}:=s_{1},$ $s_{r+1}:=t_{n}$, $t_{r+1}:=+\infty$. Indeed, the finite linear span of such functionals is dense in $L^{2}(\Omega,\mathcal{F}^{\mathcal{L}},P)$ by the tensorization property of $L^{2}$ (see proof of 2.4). But this is immediate from the result of Step 3. ∎